### v2

# Interval-Constrained Bipartite Matching over Time

Abstract
Interval-constrained online bipartite matching problem frequently occurs in medical appointment scheduling: unit-time jobs representing patients arrive online and are assigned to a time slot within their given time interval. We consider a variant of this problem where reassignments are allowed and extend it by a notion of current time, which is decoupled from the job arrival events. As jobs appear, the current point in time gradually advances. Jobs that are assigned to the current time unit become processed, which fixes part of the matching and disables these jobs or slots for reassignments in future steps. We refer to these time-dependent restrictions on reassignments as the over-time property.
We show that FirstFit with reassignments according to the shortest augmenting path rule is $\frac{2}{3}$-competitive with respect to the matching cardinality, and that the bound is tight. Interestingly, this bound holds even if the number of reassignments per job is bound by a constant. For the number of reassignments performed by the algorithm, we show that it is in $\Omega(n\log n)$ in the worst case, where $n$ is the number of patients or jobs on the online side. This result is in line with lower bounds for the number of reassignments in online bipartite matching with reassignments [14], and, similarly to this previous work, we also conjecture that this bound should be tight. Known upper bounds like the $\mathcal{O}(n\log^{2}n)$ for online bipartite matching with reassignments by Bernstein, Holm, and Rotenberg [4] do not transfer directly: while our interval constraints simplify the problem, the over-time property restricts the set of possible reassignments.
1 Introduction
Consider the process of assigning medical appointments of equal duration to patients. Patients request appointments one by one and expect an immediate answer. The time interval within which patients require an appointment is determined by their medical condition: the urgency of treatment imposes the deadline, and the amount of time needed to prepare the patient for treatment, e.g., additional diagnostics before a surgery, determines the earliest possible appointment day. As further patients arrive, previously scheduled patients with looser deadlines may be postponed in favor of urgent patients. As the time passes, a growing part of the schedule lies in the past and therefore cannot be changed, whereas appointment slots in the future can still be reassigned to different patients. The goal is to schedule as many patients as possible and to minimize changes to the schedule.
In this work, we model appointment scheduling as a variant of online bipartite matching with reassignments. In this problem, vertices of one part of the graph, here called “slots”, are known in advance, and vertices of the other part, here called “jobs”, which represent patients, are revealed one by one together with their incident edges. Upon arrival, each job is either immediately assigned to an adjacent slot or irrevocably rejected. It is possible for the algorithm to reallocate jobs to different slots in order to make place for a new job. The task is to maximize the number of assigned jobs and minimize the number of reallocations.
We extend this well-studied problem by properties motivated in the practical application. We introduce a notion of time to the online problem. As time passes, more jobs arrive and slots are processed, so that once a slot is in the past, jobs assigned to it cannot be reallocated anymore. Furthermore, our jobs feature interval constraints: a job can be allocated only to a slot between its given earliest and latest slot.
We refer to the described problem as Interval-Constrained Bipartite Matching over Time, or ICBMT for short, and formalize the problem in Section 1.3.
In the motivating medical application, slots are often assigned in a first-come-first-served manner which is then revoked in case of emergencies. This corresponds to a FirstFit algorithm that uses shortest augmenting paths (SAP) to augment the schedule when there is no feasible slot available. Aside from the practical relevance, this algorithm is a natural choice. Due to the over-time property of the problem, algorithms that leave early slots unoccupied risk missing a possible match or have to augment the matching for the current time slot. We analyze this online algorithm and give a tight bound of $\frac{2}{3}$ on the competitive ratio with respect to the cardinality of the resulting matching. Furthermore, we show that if the algorithm always selects the lexicographically smallest (lex-min) shortest path, then it incurs $\Omega(n\log n)$ reassignments in the worst case. This bound can be also adopted for the lex-max paths on a different worst-case instance.
1.1 Related work
ICBMT is a special case of the online bipartite matching problem and of the online assignment problem. In these problems, the servers (possibly with capacity greater than one) are known in advance, and clients are revealed online, each with a subset of servers the client can be served by. Typical objectives are maximizing the number of assigned clients [18] or balancing the load of the servers [5].
In the classical maximum online bipartite matching, all matching decisions are irrevocable, and the aim is to maximize the number of matched clients. No deterministic algorithm can have a competitive ratio greater than $\frac{1}{2}$, which is achieved by a greedy algorithm. Karp, Vazirani, and Vazirani present a randomized algorithm ranking, which is strictly $(1-\frac{1}{e})$-competitive, and show the tightness of the bound [18]. The proof has been later corrected, refined and generalized [13, 6, 11].
In many applications it is reasonable to allow reassignments; that is, client rejections are irrevocable, but the assigned server of an already assigned client can be changed. In this problem, it is trivial to achieve a maximum matching at the cost of a large number of reassignments. One widely studied strategy for online bipartite matching with reassignments is Shortest Augmenting Paths (SAP), which is also the basis of an optimal algorithm for offline maximum bipartite matching [17]. Grove et al. show a lower bound of $\Omega(n\log n)$ reassignments using SAP [14]. They also show a matching upper bound of $\mathcal{O}(n\log n)$ reassignments for the case of all clients having degree of two. This upper bound for SAP was later proven for arrivals in random order [9] and for trees [8]. For general graphs, Bernstein, Holm, and Rotenberg show an amortized upper bound of $\mathcal{O}(n\log^{2}n)$ reassignments for SAP [4]. Chaudhuri et al. conjecture that the lower bound is tight for general bipartite graphs [9].
Exact approaches to online bipartite matching with reassignments different from SAP have also been studied [7, 9], without improvements in the number of reassignments.
Another branch of research concerns approximation algorithms for online bipartite matching. Bosek et al. construct $(1-\epsilon)$-approximate matchings applying $\mathcal{O}((1-\epsilon)n)$ reassignments [7]. Bernstein et al. approximate the problem for every $p$-norm simultaneously. Their algorithm is $8$-competitive with $O(n)$ reassignments [5]. Gupta, Kumar, and Stein consider the assignment version of the problem, where each server can serve multiple clients, and present a $2$-competitive algorithm for maximum load minimization with at most $O(n)$ reassignments [15]. Recently, Shin et al. considered the online bipartite matching problem under a bounded number $k$ of allowed reassignments per arrival [21]. They show that a modified SAP-algorithm is $(1-\frac{2}{k+2})$-competitive for this setting and that this bound is tight already on paths.
The influence of recourse has also been studied for more general matching problems, such as minimum cost matching [20, 16] and maximum matching on general graphs [1, 3, 19].
Another closely related problem is unit job scheduling with integer release times and deadlines. The interval constraints directly correspond to release times and deadlines, and the maximum cardinality objective corresponds to throughput maximization. Here, jobs can be buffered until they are executed or the deadline comes up. Offline, this problem is optimally solved by the Earliest Deadline First strategy, which can be implemented in $O(n)$ time [12, 22]. Online, Baruah, Haritsa, and Sharma show that any algorithm that is never idle is at least 2-competitive [2]. Chrobak et al. give a barely random algorithm, an algorithm that only uses a single random bit, that is $\frac{5}{3}$-competitive [10]. Please note that in these scheduling problems, deadlines and release dates are not necessarily integer. Moreover, in contrast to ICBMT, decisions on rejections of jobs do not have to be met immediately upon job’s arrival, as jobs can remain pending for several time steps.
1.2 Our contribution
We consider a FirstFit-type algorithm that uses shortest augmenting path in case there is no available slot within a job’s interval. For this algorithm, we give a tight lower and upper bound of $\frac{2}{3}$ on the competitive ratio with respect to the cardinality of the final matching. In contrast to related problems that use recourse, the over-time nature of our problem eliminates augmenting paths that may be needed to recover an offline optimal solution to the problem.
In addition, we give an instance on which the algorithm performs $\Omega(n\log n)$ reassignments. To show this result, we have to specify the algorithm more precisely. For every new assignment there might be several shortest augmenting paths, and the choice of one impacts the result. Our algorithm always selects the lexicographically smallest shortest path, where the order is with respect to the sequence of slots in the path. This is a natural extension of the FirstFit strategy that also favors early slots for the current job. We conjecture that there is no deterministic strategy of selecting shortest paths that yields strictly better results, but a randomization over different strategies might be promising.
Our results imply that with only a constant number of reassignments for each new assignment, we retain the competitive ratio of $\frac{2}{3}$ for only linear number of reassignments in total. This shows that the difficulties of both objectives do not occur in the same instances.
Our main results are summarized in Table 1.
FirstFit $k$-FirstFit Online-EDF
assigned jobs $\frac{2}{3}$ $\frac{2}{3}$ $1$
(competitive ratio)
reassignments: LB $\Omega(n\log n)$ $\Omega(n)$ $\Omega(n^{2})$
UB ? $\mathcal{O}(n)$ $\mathcal{O}(n^{2})$
Table 1: Performance bounds for algorithms presented in this work for instances with $n$ jobs. For the case marked with ”?”, there is a known analogous bound of $\mathcal{O}(n\log^{2}n)$ for online bipartite matching [4], which does not necessarily hold for ICBMT due to its over-time aspect.
Furthermore, we consider a few special cases. If all intervals have the same fixed length, we show that FirstFit maximizes the cardinality of the matching in $\Theta(n)$ reassignments. When the number of reassignments is no concern, an online version of the Earliest Deadline First (EDF) algorithm maximizes the cardinality of the matching for $\Omega(n^{2})$ reassignments in the worst case. If the algorithms are allowed to make decisions in batches for all jobs that arrive in the same time unit, the competitive ratio does not improve for algorithms of the type that we consider. The number of reassignments in this model improves at most by a factor of $2$.
1.3 Problem definitions
In Interval-Constrained Bipartite Matching over Time, we are given processing capacity of one slot per time unit $t\in\mathbb{N}$; we identify slots with time units they belong to. Jobs $j\in\{1,\ldots,n\}$ are revealed one by one together with their arrival times $a_{j}\in\mathbb{N}_{0}$ and their intervals $[r_{j},d_{j}]$, $r_{j},d_{j}\in\mathbb{N}$, which is the set of their feasible slots. The arrival times are non-decreasing in the order of jobs’ arrival, i.e., $a_{j}\leqslant a_{j+1}$ for any $j<n$. For all jobs $j\leqslant n$, the preparation time $r_{j}-a_{j}$ is at least $1$.
Upon arrival, job $j$ has to be assigned to some slot $t\in[r_{j},d_{j}]$. Already scheduled jobs can be reassigned to different slots in order to free a slot in job $j$’s interval, but they cannot be rejected. If no reassignment combination allows to assign job $j$, the job is irrevocably rejected. Assignments to slots $t\leqslant a_{j}$ are fixed and cannot be involved in reassignments for job $j$. This models the fact that the appointment schedule for the days already past, including the current day, cannot be changed. We therefore call the slots up to $a_{j}$ fixed.
The goal is to produce online an assignment $\alpha\colon\{1,\ldots,n\}\to\mathbb{N}\cup\{\bot\}$ of jobs to slots, where $\alpha(j)=\bot$ denotes that the job is rejected. The assignment should minimize the following objectives:
  • number of rejected jobs,
  • number of reassignments.
We consider the two objectives separately.
The jobs will be denoted as tuples of their attributes $j=(a_{j},r_{j},d_{j})$, and instances as sequences of jobs.
Any instance of ICBMT has an underlying bipartite graph consisting of the job set on one side and all feasible slots on the other, with edges connecting jobs to their feasible slots. An offline augmenting path for a solution $\alpha\colon\mathcal{J}\to\mathbb{N}\cup\{\bot\}$ of ICBMT is an augmenting path with respect to the matching $\alpha$ in the underlying bipartite graph.
Remark 1.
An ICBMT solution is optimal in terms of rejections if it admits no offline augmenting paths.
Indeed, a rejection-optimal offline solution for ICBMT is a maximum matching in the underlying bipartite graph.
The main difference between ICBMT and the classical online bipartite matching with reassignments is the temporal component of ICBMT, which leads to a partially fixed assignment as time advances. However, if all jobs in an instance arrive at the same time, then ICBMT with the objective of minimizing rejections reduces to the online maximum bipartite matching. We call this special case Static ICBMT problem.
Definition 2.
Static ICBMT is a special case of the ICBMT problem, in which all jobs of an instance have the same arrival time, i.e., for any instance $I=\left((a_{j},r_{j},d_{j})\right)_{j=1}^{n}$ of Static ICBMT holds $a_{j}=a\ \text{for all}\ j\in\{1,\ldots,n\}$.
Any instance of Static ICBMT is also an instance of online bipartite matching with reassignments: indeed, since the arrival time of jobs is constant, no part of the assignment becomes fixed, and all reassignments are allowed at any time. Hence, a feasible solution for online bipartite matching is also a feasible solution for Static ICBMT on the same job set.
2 Greedy approaches
One natural and widely used appointment assignment strategy is to greedily assign the earliest possible slot to every patient. If necessary, a few other patients are reassigned.
The algorithm we call FirstFit in this work follows this greedy paradigm: it aims to assign jobs to early slots and to minimize the number of reassignments for each job arrival. It is a special case of the greedy shortest augmenting path strategy for bipartite matching that additionally considers the special structure of instances of ICBMT, namely the total order on the set of slots. The algorithm is defined as follows.
The input consists of a job $j^{\prime}=(a_{j^{\prime}},r_{j^{\prime}},d_{j^{\prime}})$ and a current slot assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$, where $\mathcal{J}\subseteq\{1,\ldots,n\}$ is the set of previously assigned jobs. Let $t^{*}$ be the earliest free slot within of after the interval of job $j^{\prime}$. FirstFit searches for a shortest sequence of reassignments which frees up one slot in the interval of $j^{\prime}$ and assigns the last job of the sequence to the slot $t^{*}$. If no such sequence of reassignments exists, job $j^{\prime}$ is rejected.
The sequence of reassignments results from a shortest augmenting path $P$ from job $j^{\prime}$ to the first free slot $t^{*}$ in a bipartite job-slot graph $G_{\alpha,j^{\prime}}$. The graph $G_{\alpha,j^{\prime}}=(V^{T}\,\dot{\cup}\,V^{J},E^{\alpha}\,\dot{\cup}\,E^{\overline{\alpha}})$ is a digraph containing a vertex for each non-fixed slot and job: $V^{T}=\left\{t\in\cup_{j\in V^{J}}[r_{j},d_{j}]\mid t>a_{j^{\prime}}\right\}$ and $V^{J}=\left\{j\in\mathcal{J}\mid\alpha(j)>a_{j^{\prime}}\right\}\,\dot{\cup}\,\{j^{\prime}\}$. There are edges $E^{\overline{\alpha}}$ from each job to its feasible slots it is not assigned to, $E^{\overline{\alpha}}=\{(j,t)\mid j\in V^{J},t\in V^{T},t\in[r_{j},d_{j}],\alpha(j)\neq t\}$, and edges $E^{\alpha}$ from each filled slot to the job assigned to it, $E^{\alpha}=\{(t,j)\mid t\in V^{T},j\in V^{J},\alpha(j)=t\}$. The path $P$ is augmenting with respect to the current assignment $\alpha$, which is interpreted as a matching.
Input: job $j^{\prime}=(a_{j^{\prime}},r_{j^{\prime}},d_{j^{\prime}})$, current slot assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$
Output: a new assignment $\alpha^{\prime}\colon\mathcal{J}\cup\{j^{\prime}\}\to\mathbb{N}\cup\{\bot\}$
$t^{*}\coloneqq\min\{t\geqslant r_{j^{\prime}}\mid\alpha^{-1}(t)=\emptyset\}$;
Find a shortest augmenting $j^{\prime}$-$t^{*}$ path $P$ in $G_{\alpha,j^{\prime}}$;
if $P$ exists then
       $\alpha^{\prime}\coloneqq\alpha\,\Delta\,\left\{(j,t)\mid(j,t)\in P\text{ or }(t,j)\in P\right\}$;
else
       $\alpha^{\prime}\coloneqq\alpha\cup(j^{\prime},\bot)$;
return $\alpha^{\prime}$
Algorithm 1 FirstFit algorithm
A shortest augmenting path can be found via breadth-first search on $G_{\alpha,j^{\prime}}$. If multiple shortest paths exist, one is selected arbitrarily. We will discuss the influence of the choice of the path on the algorithm’s performance at the end of Section 2.
Notice that FirstFit can also be used for Bipartite Matching over Time without interval constraints, where sets of feasible slots of jobs are of arbitrary structure.
2.1 Rejections by FirstFit
We start our analysis of FirstFit with the primary objective in ICBMT – the number of rejections. In general, FirstFit can make unnecessary rejections, as we will see later. On instances of Static ICBMT, however, FirstFit achieves a minimal number of rejections.
Observation 3.
FirstFit optimizes the number of rejections on instances of Static ICBMT.
Indeed, as mentioned above, Static ICBMT is a special case of maximum online bipartite matching. FirstFit, in turn, coincides with the SAP strategy for online bipartite matching complemented with a tie-breaking rule between paths of the same length: the earliest terminal slot is preferred. Hence, optimality of FirstFit follows from the optimality of SAP [9].
To prove the next optimality result, we introduce the notion of a closed interval.
Definition 4.
A closed interval in an assignment $\alpha\colon\mathcal{J}\to\mathbb{N}\cup\{\bot\}$ for ICBMT is a time interval $I\subseteq\mathbb{N}$ without empty slots such that all jobs $j\in\alpha^{-1}(I)$ have intervals $[r_{j},d_{j}]\subseteq I$.
Observe that any alternating path with respect to the matching $\alpha$ in the underlying bipartite graph that starts in a closed interval can visit only slots of this interval and jobs assigned to them.
Theorem 5.
If all jobs in an instance of ICBMT have intervals of the same length, then FirstFit minimizes the number of rejections.
Proof.
We show that for a FirstFit solution there are no offline augmenting paths in the underlying bipartite graph. To this end, we show that the interval of each rejected job lies in a closed interval, so that augmenting paths starting with the rejected job never reach an empty slot.
Consider a rejected job $j^{\prime}$ and the FirstFit assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$ upon $j^{\prime}$’s arrival. As job $j^{\prime}$ is rejected, all slots in the interval $[r_{j^{\prime}},d_{j^{\prime}}]$ are occupied.
Let $t^{+}$ be the first slot after slot $d_{j^{\prime}}$ which is still free:
$t^{+}\coloneqq\min\left\{t>d_{j^{\prime}}\mid{\alpha^{-1}(t)}=\emptyset\right\}.$
We attempt to construct an augmenting path from job $j^{\prime}$ to slot $t^{+}$. If we succeed, we find a contradiction to the rejection of $j^{\prime}$; otherwise, we show the existence of a closed interval containing the interval of $j^{\prime}$.
We construct a sequence of jobs iteratively as follows. First, denote $j_{0}\coloneqq j^{\prime}$, and let
$t_{1}\coloneqq\max\left\{d_{j}\mid j\in\alpha^{-1}([r_{j^{\prime}},d_{j^{\prime}}])\right\}\text{ and }j_{1}\coloneqq\operatorname*{arg\,max}\{d_{j}\mid j\in\alpha^{-1}([r_{j^{\prime}},d_{j^{\prime}}])\}$
be the job with the latest deadline among jobs assigned to interval $[r_{j^{\prime}},d_{j^{\prime}}]$. In the following iterations $k$ for $k\in\mathbb{N}$, we set
$t_{k}\coloneqq\max\left\{d_{j}\mid j\in\alpha^{-1}([r_{j_{k-1}},d_{j_{k-1}}])\right\}\text{ and }j_{k}\coloneqq\operatorname*{arg\,max}\left\{d_{j}\mid j\in\alpha^{-1}([r_{j_{k-1}},d_{j_{k-1}}])\right\}$
until either $t_{k}\geqslant t^{+}$ or $t_{k}=t_{k-1}$ occurs for some $k$. Note that by construction, $t_{k}\geqslant t_{k-1}$ is true for all $k\in\mathbb{N}$.
If the process terminates with $t_{k}\geqslant t^{+}$, then the sequence $Q\coloneqq(j^{\prime},\alpha(j_{1}),j_{1},\alpha(j_{2}),j_{2},\ldots,j_{k},t^{+})$ is an augmenting path for job $j^{\prime}$ and assignment $\alpha$. Indeed, slot $\alpha(j_{\ell})$ is feasible for job $j_{\ell-1}$ by construction of the job sequence, as $\alpha(j_{\ell})\in[r_{j_{\ell-1}},d_{j_{\ell-1}}]$ for all $\ell\in\{1,\ldots,k-1\}$. The free slot $t^{+}$ is feasible for the last job of the sequence $j_{k}$ due to the following relations:
$r_{j_{k}}\leqslant\alpha(j_{k})\leqslant d_{j_{k-1}}=t_{k-1}<t^{+}.$
Finally, slots in $Q$ are not fixed, since $\alpha(j_{\ell})\geqslant r_{j_{\ell-1}}\geqslant r_{j^{\prime}}$ is true by construction for all $\ell\in\{1,\ldots,k\}$. Although $Q$ is not necessarily the shortest augmenting path used by FirstFit, the existence of sequence $Q$ guarantees the existence of a shortest such path. Hence, job $j^{\prime}$ would not be rejected by FirstFit.
So suppose the sequence construction stops at iteration $k$ with $t_{k}=t_{k-1}$ and $t_{k}<t^{+}$. Here, we show that there exists a closed interval for $\alpha$ containing the interval of $j^{\prime}$. For all jobs $j\in\alpha^{-1}([r_{j_{k-1}},d_{j_{k-1}}])$ we have $d_{j}\leqslant t_{k}$; consequently, since all jobs have intervals of the same length, for any job $j$ with $\alpha(j)<r_{j_{k-1}}$ it also follows that $d_{j}<d_{j_{k-1}}=t_{k}$. Hence, we set $t_{k}$ as the right bound of the sought closed interval.
To determine the left bound of the closed interval, we consider the latest empty slot before job $j^{\prime}$’s interval:
$t^{-}\coloneqq\max\left\{t<r_{j^{\prime}}\mid{\alpha^{-1}(t)}=\emptyset\right\}.$
Observe that for any job $j$ assigned to a slot $\alpha(j)>t^{-}$, slot $t^{-}$ cannot be feasible – otherwise, job $j$ would have been assigned by FirstFit to the earlier slot $t^{-}$. Hence, for any job $j$,
$\alpha(j)>t^{-}\text{ implies }r_{j}>t^{-}.$
Combining the two bounds, we obtain
$r_{j}\geqslant t^{-}+1\quad\text{and}\quad d_{j}\leqslant t_{k}\quad\text{for all}\ j\in\alpha^{-1}([t^{-}+1,t_{k}]),$
that is, every slot of the interval $I\coloneqq[t^{-}+1,t_{k}]$ is occupied by a job that requires a slot in this time interval.
No alternating job-slot sequence with respect to $\alpha$ starting with a slot in $I$ can ever leave interval $I$. Therefore, jobs assigned to $I$ in $\alpha$ will not be replaced by jobs arriving later. Thus, also in the final FirstFit solution no offline alternating path starting in $I$, or particularly in $[r_{j^{\prime}},d_{j^{\prime}}]\subseteq I$, leads to an empty slot. ∎
Theorem 6.
FirstFit is at most $\frac{2}{3}$-competitive for ICBMT w.r.t. rejections when jobs have at least two different interval lengths.
Proof.
We define two job types: urgent jobs with preparation time of $1$ and interval length $\delta$, and regular jobs with preparation time $\Omega>1$ and intervals of length $\Delta\geqslant 3\delta-2$. Consider now the following instance $\mathcal{I}$ of ICBMT (see Figure 1). First, $\ell\coloneqq\delta-1$ regular jobs $j_{i}$ for $i\in\{1,\ldots,\ell\}$ arrive at time $1-\Omega$ (we extend the time scale to negative numbers to simplify the notation). Their interval is $[1,\Delta]$, and FirstFit assigns them to slots $\alpha(i)=i$ for $i\in\{1,\ldots,\ell\}$. Next, for each $i\in\{1,\ldots,\delta\}$, an urgent job $u_{i}$ with interval $[i,i+\ell]$ is revealed at time $i-1$. Upon arrival of job $u_{i}$, the first $\ell+i-1$ slots are occupied, so $u_{i}$ is assigned to slot $\ell+i$, the last slot of its interval.
Finally, further $\ell$ urgent jobs $v_{i}$, $i\in\{1,\ldots,\ell\}$ with interval $[\delta,\delta+\ell]$ are revealed at time $\ell$. All slots in their interval are occupied. The assignment up to the current time point $\ell$ cannot be modified, so the regular jobs cannot be reassigned.
$t$$\ell$$\delta$$\Delta$$\ell+\delta$$\ell=\delta-1$$\delta$$\ell$
Figure 1: Upper bound instance for $\delta=4$, $\ell=3$. The arrival order of jobs is from top to bottom. The arrival times are marked with dots and the FirstFit assignment with yellow circles. The grayed-out part of the solution is fixed upon arrival of the last $\delta$ jobs. The dashed jobs are rejected.
Hence, jobs $v_{i}$, $i\in\{1,\ldots,\ell\}$, are all rejected by FirstFit, so FirstFit achieves the objective value $\text{FF}(\mathcal{I})=2\delta-1$. An optimal assignment $\beta$ assigns all jobs to a feasible slot:
$\begin{array}[]{ll}\beta(j_{i})\,=\delta+\ell+i,&1\leqslant i\leqslant\ell,\\ \beta(u_{i})=i,&1\leqslant i\leqslant\delta,\\ \beta(v_{i})=\delta+i,&1\leqslant i\leqslant\ell;\end{array}$
hence, $\text{OPT}(\mathcal{I})=3\delta-2$. From
$\frac{\text{FF}(\mathcal{I})}{\text{OPT}(\mathcal{I})}=\frac{2\delta-1}{3\delta-2}\underset{\delta\to\infty}{\longrightarrow}\frac{2}{3}$
follows that the competitive ratio of FirstFit is not greater than $\frac{2}{3}$. ∎
We complement this result with a tight lower bound. Even if we drop the interval constraints and allow for arbitrary sets of feasible slots, FirstFit is at least $\frac{2}{3}$-competitive.
Theorem 7.
FirstFit is at least $\frac{2}{3}$-competitive for ICBMT with respect to the number of rejections.
Proof.
Consider an arbitrary instance of ICBMT. Let $\mathcal{J}_{\text{FF}}$ be the set of jobs assigned by FirstFit, and $\mathcal{J}_{\text{OPT}}$ be the set of jobs assigned in an optimal solution; let $f\colon\mathcal{J}_{\text{FF}}\to\mathbb{N}$ and $o\colon\mathcal{J}_{\text{OPT}}\to\mathbb{N}$ be the corresponding assignments. To compare the cardinalities of the two assignments, we construct alternating job-slot sequences that help to partition the job set. For each job $j_{0}\in\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}$ rejected by FirstFit but assigned by the optimal algorithm, we construct a sequence $Q(j_{0})$ by first adding $j_{0}$ and its optimal slot $s_{0}=o(j_{0})$ to it. Next, in each iteration $k\in\mathbb{N}$, we consider the job $j_{k}=f^{-1}(s_{k-1})$ that is assigned by FirstFit to the optimal slot of the previous job. Job $j_{k}$ and its optimal slot $s_{k}=o(j_{k})$ are added to the sequence $Q(j_{0})$. We repeat these iterations until we reach a slot $s_{k}$ that is free in $f$ or a job $j_{k}$ that is not assigned by $o$.
Our construction yields two types of sequences. In each sequence, the first job is in $\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}$, the intermediate jobs are in $\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}$, and the last job is either in $\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}$ as well, if the sequence ends with a slot, or in $\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}$. We call the sequences that end with a job in $\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}$ and a slot even and those that end in $\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}$ odd.
We show that each even sequence contains at least three jobs. If a sequence $Q(j_{0})$ for some rejected job $j_{0}$ contains only one job, then $Q(j_{0})=(j_{0},o(j_{0}))$, and slot $o(j_{0})$ is free in assignment $f$. So FirstFit would have assigned $j_{0}$ to this slot.
Now consider an even sequence with two jobs $Q(j_{0})=(j_{0},s_{0},j_{1},s_{1})$, where $s_{0}=o(j_{0})$, $s_{1}=o(j_{1})$ and $f(j_{1})=s_{0}$ (see Fig. 2).
$j_{0}$$j_{1}$$s_{0}$$s_{1}$
Figure 2: An even alternating sequence with two jobs.
Since slot $s_{1}$ is free in the FirstFit solution and job $j_{0}$ is rejected, slot $s_{1}$ is infeasible for $j_{0}$. Furthermore, $j_{1}$ arrived before $j_{0}$, and $s_{0}<s_{1}$, because FirstFit assigned $s_{0}$ to $j_{1}$. Upon arrival of $j_{0}$, FirstFit would thus reassign $j_{1}$ instead of rejecting job $j_{0}$, which is feasible because it yields the optimal assignment $o$.
Next, we express the number of jobs assigned or rejected by both algorithms through the number of odd and even sequences $n_{\text{odd}}$ and $n_{\text{even}}$, respectively. For the odd sequences, we use the trivial bound $n_{\text{odd}}\leqslant\left|\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}\right|$. For the even sequences, we have shown that each sequence contains at least three jobs. This yields a bound of $n_{\text{even}}\leqslant\frac{1}{2}\left|\mathcal{J}_{\text{FF}}\cap\mathcal{J}_{\text{OPT}}\right|$. As jobs in $\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}$ only occur at the start of a sequence, we have
$\left|\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}\right|=n_{\text{odd}}+n_{\text{even}}\leqslant\left|\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}\right|+\frac{1}{2}\left|\mathcal{J}_{\text{FF}}\cap\mathcal{J}_{\text{OPT}}\right|\;.$
Finally, we obtain
$\left|\mathcal{J}_{\text{OPT}}\right|$ $=\left|\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}\right|+\left|\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}\right|\leqslant\frac{3}{2}\left|\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}\right|+\left|\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}\right|\leqslant\frac{3}{2}\left|\mathcal{J}_{\text{FF}}\right|\;.$
2.2 Reassignments by FirstFit
Another natural performance measure for online algorithms for ICBMT is the number of reassignments made by the algorithm. Before we present our analysis of FirstFit for this measure, note that we did not specify how ties between shortest augmenting paths from the current job to the earliest free slot are resolved. Although each of the shortest paths yields the same number of reassignments in that iteration, we will see that the resulting solutions lead to different numbers of reassignments in later iterations.
All possible shortest augmenting paths can be uniquely represented and compared by the sequence of slots they visit on the way from the new job to the target slot. One possible rule for path selection is to choose the lexicographically smallest sequence of slots. Intuitively, this rule favors early slots for incoming jobs, and the path can be found in polynomial time.
We show in Lemma 8 that with this path choice rule, which we denote lex-min for short, FirstFit makes at least $\mathcal{O}(n\log n)$ reassignments in the worst case. Observe that this bound coincides with the lower bound on the worst-case number of reassignments for the SAP strategy for online bipartite matching [14].
Lemma 8.
In the worst case, FirstFit makes $\Omega(n\log n)$ reassignments on instances with $n$ jobs, even on instances of Static ICBMT.
Proof.
Here, we provide only a sketch of the proof; for the full proof, see Appendix A.
Consider an upper-triangle instance with $N=2^{n}$ jobs arriving at time 00. Jobs are enumerated in the order of arrival, job $i\in\{1,\ldots,N\}$ has interval $[1,N+1-i]$, see Figure 3. To describe how FirstFit works on this instance, we divide the set of jobs into $n+1$ subsequent phases: phase $k\leqslant n$ contains $\frac{N}{2^{k}}$ jobs, and phase $n+1$ contains one job.
In the first phase, the first half of all the jobs arrives and is assigned by FirstFit to the first $\frac{N}{2}$ slots. In the second phase, the half of the remaining jobs arrives. As there is no free slot for any of these new jobs, each new job causes a reassignment of one of the jobs of phase one. According to the lex-min rule for choosing augmenting paths, jobs of phase $2$ are assigned to the first $\frac{N}{4}$ slots, and the first $\frac{N}{4}$ jobs of phase $1$ are reassigned to the third block of $\frac{N}{4}$ slots, i.e., to slots $\frac{N}{2}+1,\ldots,\frac{3N}{4}$. In each subsequent phase $k$, the half of already present jobs gets reassigned. The chain of reassignments caused by job $i$ ends with slot $i$, and each job moves $\frac{N}{2^{k-1}}$ slots to the right, so each chain contains $12^{k-1}-$ 1 reassignments (see Fig. 3). Hence, in phase $k$ there are $\frac{N}{2^{k}}(2^{k-1}-1)=\frac{N}{2}-\frac{N}{2^{k}}$ reassignments in total. For the derivation of these calculations see Appendix A.
$t$Phase 1Phase 2Phase 3
Figure 3: Example instance with the reassignments in phase $3$. Under lex-min rule, jobs of the current phase occupy the first slots in their order of arrival, and all other jobs are assigned in parallel diagonal blocks. The reassignment chains propagate this block structure through the instance.
Hence, the total number of reassignments up to phase $\log(N)$ is
$\sum_{k=2}^{\log(N)}\frac{N}{2}-\frac{N}{2^{k}}$ $=\frac{N}{2}(\log(N)-1)-N\sum_{k=2}^{\log(N)}\frac{1}{2^{k}}$
$=\frac{N}{2}(\log(N)-1)-N\left(\frac{1}{2}-\frac{1}{N}\right)\in\Omega(N\log(N))\;.$
Observe that on the ICBMT instance in Lemma 8, a different path choice rule, e.g., the lex-max rule that chooses the lexicographically maximal sequence of visited slots, would lead to only $\frac{N}{2}$ reassignments in total.
However, the lex-max rule is subobtimal too: consider an instance similar to that in Figure 3, but in which the second half of the jobs have right-aligned intervals $[i-\frac{N}{2},\frac{N}{2}]$ for $i\in\{\frac{N}{2}+1,\ldots,N\}$. By reasoning analogous to the proof of Lemma 8, the lex-max path choice rule yields $\mathcal{O}(n\log n)$ reassignments on this instance, whereas lex-min requires only $\frac{N}{2}$ reassignments.
In general, we were not able to identify any dominant rule for the choice of paths. Hence, we conjecture the following.
Conjecture 9.
There is no deterministic rule for the choice of paths for FirstFit that dominates all other possible rules on all instances of ICBMT.
For online bipartite matching, there is a known upper bound of $\mathcal{O}(n\log^{2}n)$ reassignments under the Shortest Augmenting Path strategy. This bound holds in particular for Static ICBMT as a special case of online bipartite matching, but not necessarily for ICBMT in general. Unfortunately, we were not able to show an analogous upper bound for ICBMT. Our problem differs from online bipartite matching in two aspects: the interval structure of feasible sets and the temporal component that makes parts of the graph become fixed. While the interval constraints possibly reduce the complexity of the problem by restricting possible adversarial instances, the temporal aspect of ICBMT instances makes any SAP-based algorithm less potent, since there are fewer augmenting paths available, as Figure 4 demonstrates. We conjecture that the worst-case number of reassignments for FirstFit on ICBMT is not smaller than that for reassignments on online bipartite matching.
$t$$n-4$
Figure 4: An instance with $n=8$ jobs with the FirstFit assignment, identical to an SAP assignment, upon arrival of the last job. The reassignments by FirstFit in ICBMT are shown in red. Online bipartite matching admits a shorter augmenting path of length $2$, shown in blue dashed.
Finally, we again consider a special case of instances with uniform interval length, assuming that these instances allow for better bounds also with respect to the number of reassignments.
Lemma 10.
If all $n$ jobs have the same interval length $\delta$, then FirstFit makes $\mathcal{O}(\delta n)$ reassignments.
Proof.
We first show that in the described instances, FirstFit never reassigns a job to an earlier slot. To begin with, the properties of FirstFit imply that the target slot at the end of any reassignment sequence is later than the start slot of the sequence, which lies within the new job’s interval. Hence, any reassignment sequence of FirstFit ends with a forward reassignment – a reassignment to a later slot.
Next, suppose that some reassignment sequence with respect to the start assignment $\alpha$ contains a backward reassignment
$j\to t\coloneqq\alpha^{-1}(j^{\prime})\quad\text{with }t<\alpha(j),$
followed by a forward reassignment $j^{\prime}\to t^{\prime}$, i.e., $t^{\prime}>t$.
If $t^{\prime}<\alpha(j)$ and thus $t^{\prime}\in[t,\alpha(j)]$, then slot $t^{\prime}$ is feasible for job $j$, so the reassignment sequence can be shortened by assigning $j\to t^{\prime}$ directly. If $t^{\prime}>\alpha(j)$, then the slots $t$ and $t^{\prime}$ are less than one interval length apart, since they are both feasible for job $j^{\prime}$. Now consider the job $j_{0}$ that precedes $j$ in the sequence and is reassigned to the slot $\alpha(j)$ (job $j_{0}$ can also be the new job that initiates the reassignments). Since its interval contains slot $\alpha(j)$, it also contains at least one of the slots $t$ and $t^{\prime}$. In both cases, the reassignment sequence can be shortened by assigning job $j_{0}$ directly to the slot $t$ or $t^{\prime}$.
Consequently, FirstFit makes only forward reassignments on instances with uniform interval length $\delta$. Then, every job can be reassigned at most $\delta-1$ times, yielding the overall upper bound of $(\delta-1)n=\mathcal{O}(\delta n)$ reassignments for instances with $n$ jobs. ∎
If the interval length is considered fixed, we obtain a bound of $\mathcal{O}(n)$ reassignments. We show that this bound is tight.
Lemma 11.
On instances of ICBMT with uniform interval length, FirstFit makes $\Omega(n)$ reassignments in the worst case.
Proof.
Consider the following instance with interval length $\delta$ (see Fig. 5). First, $\delta-1$ identical jobs with intervals $[1,\delta]$ arrive and are assigned to the first $\delta-1$ slots. Next, $k$ jobs with intervals $[i+1,i+\delta]$ for $i\in\{1,\ldots,k\}$ arrive and become assigned each to its one but last slot. The first $k+\delta-1$ slots are now occupied. Now one further job with interval $[1,\delta]$ arrives. Each of the $k$ jobs with later intervals is shifted by one slot forward, yielding $k$ reassignments for an instance with $k+\delta$ jobs.
[…]
Figure 5: An example instance at the moment of arrival of the last job. Yellow circles show the assignment at that moment, arrows indicate caused reassignments.
2.3 Limited number of reassignments
Next, we introduce a modification of FirstFit which makes a limited number of reassignments per job arrival, analogously to the SAP strategy considered in [21]. FirstFit with path limit $k$, or $k$-FirstFit for short (Alg. 2), is only allowed to use augmenting paths whose length does not exceed $12k+$ 1 for a fixed $k\in\mathbb{N}$, that is, paths which contain at most $k$ already scheduled jobs. Thus, k-FirstFit reassigns at most a constant number of jobs per new assignment.
Input: job $j^{\prime}=(a_{j^{\prime}},r_{j^{\prime}},d_{j^{\prime}})$, current slot assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$
Output: new assignment $\alpha^{\prime}\colon\mathcal{J}\cup\{j^{\prime}\}\to\mathbb{N}\cup\{\bot\}$
$t^{*}\coloneqq\min\{t\geqslant r_{j^{\prime}}\mid\alpha^{-1}(t)=\emptyset\}$;
Find a shortest augmenting $j^{\prime}$-$t^{*}$ path $P$ in $G_{\alpha,j^{\prime}}$;
if $\left|P\right|\leqslant 2k+1$ then
       $\alpha^{\prime}\coloneqq\alpha\,\Delta\,\left\{(j,t)\mid(j,t)\in P\text{ or }(t,j)\in P\right\}$;
else
       $\alpha^{\prime}\coloneqq\alpha\cup(j^{\prime},\bot)$;
return $\alpha^{\prime}$
Algorithm 2 FirstFit with path limit $k$ ($k$-FirstFit)
Clearly, $k$-FirstFit makes at most $kn=\mathcal{O}(n)$ reassignments. Interestingly, both Lemma 7 and the example in Lemma 6 also hold for FirstFit with path limit. Hence, we obtain an assignment strategy which is bounded in both the number of rejections and reassignments.
Corollary 12.
$k$-FirstFit assignment strategy for ICBMT is $\frac{2}{3}$-competitive with respect to the number of rejections and makes at most $k\cdot n$ reassignments.
3 Proactive reassignments
So far, we have assumed that reassignments are made only if it is otherwise impossible to assign the job just arrived. Now we investigate an assignment algorithm that strategically reassigns jobs upon arrival of new requests.
The Online-EDF algorithm (Alg. 3) mimics the optimal offline assignment strategy Earliest Deadline First (EDF). It maintains a maximal matching in the future part of the schedule and rejects jobs that cannot be matched in addition to the already scheduled ones. Ties in the ordering of jobs in the schedule are broken according to non-decreasing deadlines.
Input: job $j^{\prime}=(a_{j^{\prime}},r_{j^{\prime}},d_{j^{\prime}})$, current slot assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$
Output: new assignment $\alpha^{\prime}\colon\mathcal{J}\cup\{j^{\prime}\}\to\mathbb{N}\cup\{\bot\}$
$S\coloneqq\{j^{\prime}\}\cup\{j\in\mathcal{J}\mid\alpha(j)>a_{j^{\prime}}\text{ and }d_{j}>d_{j^{\prime}}\}$;
Sort $S$ by deadlines;
foreach $j\in S$ do
       if no free slot in $[r_{j},d_{j}]$ then
             $\alpha^{\prime}\coloneqq\alpha\cup(j^{\prime},\bot)$;
              /* reject $j^{\prime}$/
             return $\alpha^{\prime}$
      else
             $\alpha^{\prime}(j)\coloneqq\min\{t\in[r_{j},d_{j}]\mid(\alpha^{\prime})^{-1}(t)=\emptyset\}$;
              /* choose the first free slot */
return $\alpha^{\prime}$
Algorithm 3 Online-EDF
Observe that, in contrast to scheduling problems, where jobs with earlier deadlines are prioritized and may push out other jobs from the schedule, Online-EDF for ICBMT respects the arrival order of jobs and rejects later jobs with an earlier deadline, if necessary.
Theorem 13.
Online-EDF minimizes the number of rejections.
Proof.
We show that a final Online-EDF solution admits no offline augmenting paths. To this end, we show for each rejected job that its interval is contained in a closed interval.
Let $j^{\prime}$ be a job rejected by Online-EDF, and let $\alpha\colon\mathcal{J}\to\mathbb{N}$ be the slot assignment upon its arrival. We first show that a closed interval for assignment $\alpha$ existed at that time point, and argue afterwards that this interval remains in the final solution.
As job $j^{\prime}$ was rejected, some job $z\in\mathcal{J}$ did not obtain a slot during the reassignments triggered by $j^{\prime}$. Let $\beta\colon\overline{\mathcal{J}}\to\mathbb{N}$, where $\overline{\mathcal{J}}\subseteq\mathcal{J}\cup\{j^{\prime}\}$, denote the auxiliary slot assignment at the iteration in which $z$ is considered: $\beta$ is identical to $\alpha$ on the slots up to the current slot $t_{0}\coloneqq a_{j^{\prime}}$ and on jobs with deadlines up to $d_{j^{\prime}}$, and assigns the remaining jobs of set $\overline{\mathcal{J}}$ according to the EDF rule. Note that either $j^{\prime}=z$ or $d_{j^{\prime}}<d_{z}$ is true; that is, the newly arrived job either got no slot or has pushed out some other job with a later deadline. We show that there is a closed interval for assignment $\beta$ that contains the interval $[r_{z},d_{z}]$ of the pushed-out job. To find the left interval bound, we proceed similarly to the proof of Theorem 5: we iteratively construct a closure of the interval $[l,u]\coloneqq[r_{z},d_{z}]$. In each iteration, we extend the interval towards the earliest release time of the jobs assigned to the current interval. Formally, let $r_{0}\coloneqq r_{z}$. For $i\in\mathbb{N}$ define
$r_{i}\coloneqq\min\{r_{j}\mid j\in\beta^{-1}([r_{i-1},u])\}\qquad\text{and}\qquad j_{i}\coloneqq\operatorname*{arg\,min}\{r_{j}\mid j\in\beta^{-1}([r_{i-1},u])\}.$
Observe that $0\leqslant r_{i}\leqslant r_{i-1}$ for all $i\in\mathbb{N}$, that is, the sequence is monotonously decreasing and bounded from below. Hence, there exists $k\in\mathbb{N}$ for which $r_{k+1}=r_{k}$ occurs for the first time. By construction, for each job $j\in\beta^{-1}\left([r_{k},u]\right)$ holds that $r_{j}\geqslant r_{k+1}=r_{k}$. We now show that for these jobs also $d_{j}\leqslant u$ is true. For any job $j$ assigned by $\beta$, the EDF order implies the following: for any further job $y$ with $\beta(y)\in\big{[}r_{j},\beta(j)\big{)}$ we have $d_{y}\leqslant d_{j}$. Otherwise, $j$ would have been assigned to the earlier slot $\beta(y)$, since, at the moment when $y$ got assigned, both jobs $j$ and $y$ were present and not yet fixed.
We use this auxiliary statement to prove the upper bound on the deadlines of jobs assigned to the interval $[r_{k},u]$ by induction. For any $j\in\beta^{-1}([l,u])$ we have $d_{j}\leqslant u$ due to the EDF order. Now, we show that if $d_{j}\leqslant u$ holds for any job $j$ in $\beta^{-1}\left([r_{i-1},u]\right)$ for some $i\in\mathbb{N}$, then it also holds for all jobs in $\beta^{-1}\left([r_{i},u]\right)$. Consider job $j_{i}$ with release time $r_{i}$. Since $\beta(j_{i})\in[r_{i-1},u]$, we have $d_{j_{i}}\leqslant u$. For all $j$ with $\beta(j)\in\beta^{-1}\left[r_{i},r_{i-1}\right]\subseteq\big{[}r_{i},\beta(j_{i})\big{)}$ we have $d_{j}\leqslant d_{j_{i}}\leqslant u$ by the auxiliary statement above. Since the induction hypothesis holds for $[r_{i-1},u]$, we conclude that the hypothesis holds for the entire interval $[r_{i},u]$. By induction, any job assigned by Online-EDF to the interval $[r_{k},u]$ has a deadline $d_{j}\leqslant u$. To sum up, any job assigned to the interval $I\coloneqq[r_{k},u]$ during the reassignment action of job $j^{\prime}$ has $r_{j}\geqslant r_{k}$ and $d_{j}\leqslant u$. To show that interval $I$ is closed under $\beta$, it remains to show that it is fully occupied. Assume that there is an unoccupied slot $t_{u}\in[r_{k},u]$; let $t_{u}$ be the latest such slot. Clearly, $t_{u}<r_{z}=r_{0}$. Since the sequence $(r_{i})_{i=0}^{k}$ is strictly monotonously decreasing, there exists a unique $i\in\mathbb{N}$ such that $t_{u}\in[r_{i},r_{i-1})$. Then consider job $j_{i}$. The empty slot $t_{u}$ is feasible for $j_{i}$, and $j_{i}$ is assigned to a later slot $\beta(j_{i})\geqslant r_{i-1}>t_{u}$, which contradicts the Online-EDF assignment rule. Consequently, all slots in the interval $[r_{k},u]$ are occupied.
Hence, interval $I$ is closed under the auxiliary assignment $\beta$. Next, we show that is it also closed for assignment $\alpha$. If $j^{\prime}=z$, then $\alpha^{-1}(I)=\beta^{-1}(I)$, as all jobs assigned by $\beta$, except for $j^{\prime}$, are also part of assignment $\alpha$ and can be assigned only to slots in $I$. If $j^{\prime}\neq z$, then $\alpha^{-1}(I)=\beta^{-1}(I)\setminus\{j^{\prime}\}\cup\{z\}$, and interval $I$ is still closed under $\alpha$, since $[r_{z},d_{z}]\subseteq I$. Moreover, we have $\beta(j^{\prime})\in I$ (otherwise $\alpha$ would be assigning $\left|I\right|+1$ jobs to interval $I$), and hence $[r_{j^{\prime}},d_{j^{\prime}}]\subseteq I$.
It remains to show that interval $I$ stays closed in the final Online-EDF assignment. Observe that Online-EDF makes only forward reassignments. None of the jobs assigned to $I$ upon arrival of $j^{\prime}$ can be moved forward outside $I$ due to their deadlines. Therefore, none of the jobs arriving after $j^{\prime}$ will be able to push out some job from $I$. Hence, the same set of jobs remains assigned to interval $I$ in the final solution, and so $I$ is again closed. Since the interval of the rejected job $j^{\prime}$ lies in a closed interval, there is no augmenting path for the final matching connecting $j^{\prime}$ and an empty slot. ∎
Clearly, Online-EDF is inefficient in terms of the number of reassignments. We show that it achieves the trivial upper bound of $\Omega(n^{2})$ reassignments per $n$ jobs.
Lemma 14.
In the worst case, Online-EDF makes $\Omega(n^{2})$ reassignments, even on instances of Static ICBMT.
Proof.
Consider again an instance consisting of $n$ jobs with monotonously decreasing deadlines, as in the proof of Lemma 7 (see Fig. 3). Upon each job arrival, all already present jobs will be reassigned one slot further, in order to assign the new job to the first slot. Hence, the algorithm makes $\sum_{i=1}^{n}(i-1)=\frac{n(n-1)}{2}$ reassignments in total. ∎
4 Batching per time unit
In the problem definition in the introduction, we assume that jobs are processed strictly in the order of their arrival. Now we define an extended version of ICBMT, in which we are allowed to buffer the jobs arriving during one time unit and to assign them all at once, thus obtaining potentially better solutions.
Definition 15 (ICBMT with Batching, ICBMT-B).
In ICBMT with Batching, the assignment is recomputed once per time unit. That is, all jobs with the same arrival time are buffered and assigned simultaneously.
We call algorithms for ICBMT-B batching algorithms. They additionally feature a deterministic behavior on input consisting of batches, i.e., sets of jobs. We say that a non-batching algorithm $\mathcal{A}$ and a batching algorithm $\mathcal{A}^{\prime}$ correspond to each other, if their behavior on single jobs is identical.
Lemma 16.
Let $\mathcal{A}$ be an algorithm for ICBMT, and let $\mathcal{A}^{\prime}$ be a corresponding batching algorithm. For any instance $\mathcal{I}$ with $n$ jobs, there exists an instance $\mathcal{I}^{\prime}$ with less than $2n$ jobs such that algorithm $\mathcal{A}^{\prime}$ makes the same number of rejections and reassignments on instance $\mathcal{I}$ as algorithm $\mathcal{A}$ on instance $\mathcal{I}$.
Proof.
We transform instance $\mathcal{I}$ to $\mathcal{I}^{\prime}$ as follows. For each time unit $t$ with $k>1$ arrivals, we split slot $t$ to $k$ consequent slots $t^{(i)}$, $i\in\{1,\ldots,k\}$. The jobs arriving at time $t$ now receive pairwise different arrival times $t^{(i)}$ in the order of their arrival. Next, we add $k-1$ dummy jobs $\left(-1,t^{(2)},t^{(k)}\right)$ to the instance. The total number of dummy jobs is less than $n$. They block all the additionally created slots already at the start of the algorithm’s execution, so that starting from time 00, only original slots and jobs remain in the instance. As the original jobs have pairwise different arrival times, algorithm $\mathcal{A}^{\prime}$ processes them identically to the corresponding non-batching algorithm $\mathcal{A}$. ∎
Hence, batching algorithms make amortized at least the half of the rejections and reassignments per job arrival compared to their counterparts without batching. The competitive ratio of a batching algorithm with respect to rejections is thus at most $\frac{c+1}{2}$, where $c$ is the competitive ratio of the corresponding non-batching algorithm.
For the algorithms considered in this work, however, we show a stronger result: batching does not improve the competitive ratio.
Lemma 17.
The competitive ratios of Online-EDF and FirstFit with respect to rejections cannot be improved by batching.
Proof.
Online-EDF with batching cannot make fewer rejections than the already optimal Online-EDF. As for FirstFit, we slightly modify the worst-case instance family described in Theorem 7 (see Fig. 2) by assigning pairwise disjoint arrival times to the long jobs. ∎
Moreover, Lemma 16 shows that batching has no effect on our asymptotic results on numbers of reassignments.
Corollary 18.
Batching does not improve the worst-case asymptotic number of reassignments of any algorithm for ICBMT.
Acknowledgments
We thank Christina Büsing for support and helpful discussions. This work was partially supported by DFG (German Research Foundation) – RTG 2236/2.
Appendix A Proof of Lemma 8
Lemma 19.
In the worst case, FirstFit makes $\Omega(n\log n)$ reassignments on instances with $n$ jobs, even on instances of Static ICBMT.
Proof.
Consider an upper-triangle instance with $N=2^{n}$ jobs arriving at time 00. Jobs are enumerated in the order of arrival, job $i\in\{1,\ldots,N\}$ has time window $[1,N+1-i]$, see Fig. 3.
We consider only the time interval $[1,N]$, since all job time windows are included in this interval. We partition the job set into $n+1$ phases: phase $k\leqslant n$ contains $\frac{N}{2^{k}}=2^{n-k}$ jobs with indices $N-\frac{N}{2^{k-1}}+1,\ \ldots,\ N-\frac{N}{2^{k}}$, and phase $n+1$ contains one last job.
Claim 1.
When processing jobs of phase $k$ for $k\leqslant n$, FirstFit makes $\frac{N}{2^{k}}(2^{k-1}-1)$ reassignments, and at the end of phase $k$, the assignment $\alpha_{k}$ has the following structure. Jobs arrived so far are split into $12^{k}-$ 1 blocks of $\frac{N}{2^{k}}$ subsequent jobs each; block $b\in\mathbb{N}$ contains jobs $\frac{N}{2^{k}}(b-1)+1,\ \ldots,\ \frac{N}{2^{k}}b$. The time horizon $[1,N]$ of the instance is also split into $2^{k}$ intervals of length $\frac{N}{2^{k}}$. Let $L_{k}\coloneqq\frac{N}{2^{k}}$ denote the block size at the end of phase $k$. Jobs of block $b\leqslant 2^{k}-1$ are assigned to the $(b+1)$-th last interval in the order of arrival, i.e.,
$\alpha_{k}\colon j_{(b,i)}^{k}\mapsto L_{k}(2^{k}-b-1)+i,$
where $j_{(b,i)}^{k}=L_{k}(b-1)+i$ is the $i$-th job of the $b$-th block in phase $k$, $11\leqslant j\leqslant 2^{k}-$ 1, $1\leqslant i\leqslant L_{k}$.
$\triangleright$ Proof by induction.
$k=1$: Clearly, the first $\frac{N}{2}$ jobs are assigned by FirstFit to the slots $[1,\frac{N}{2}]$ in the order of their arrival, i.e.,
$\alpha_{1}\colon[1,\frac{N}{2}]\to[1,\frac{N}{2}],\quad i\mapsto i.$
This corresponds to one block of jobs, assigned to the second last (that is, first) interval of length $\frac{N}{2}$. No reassignments are necessary.
$k-1\to k$: Assume that upon arrival of jobs of phase $k$, the slot assignment $\alpha_{k-1}$ has the structure described above; in particular, jobs of block
$B_{b}\coloneqq\left\{L_{k-1}(b-1)+1,\ldots,L_{k-1}(b-1)\right\}$
are assigned to the interval
$I_{b}\coloneqq\left[L_{k-1}(2^{k-1}-b-1)+1,\ L_{k-1}(2^{k-1}-b)\right],$
where $b\in\{1,\ldots,2^{k-1}-1\}$. We denote by $I_{0}\coloneqq[L_{k-1}(2^{k-1}-1),N]$ the rightmost interval, which contains free slots. Observe that the intervals are enumerated from right to left. First, we show that for any job of phase $k$, at least $12^{k-1}-$ 1 reassignments are necessary. jobs of phase $k$ have deadlines $\frac{N}{2^{k-1}}$ or less. Hence, they require slots within the leftmost interval $I_{2^{k-1}-1}$, which is completely occupied by jobs of block $B_{2^{k-1}-1}$. For any $b\in\{1,\ldots,2^{k-1}-1\}$, jobs of block $B_{b}$ have a deadline not greater than the deadline of job $j_{(b,1)}$, which is
$N+1-L_{k-1}(b-1)-1=L_{k-1}(2^{k-1}-b+1),$
which is the right end point of the interval $I_{b-1}$. Hence, for any $b\in\{1,\ldots,2^{k-1}-1\}$, jobs of block $B_{b}$ can be reassigned to interval $I_{b-1}$ the latest, replacing jobs of block $B_{b-1}$. Consequently, any reassignment path from the interval $I_{2^{k-1}-1}$, where the time windows of new jobs are located, to the empty slots in the interval $I_{0}$ will pass through all intermediate blocks and intervals. That is, at least $12^{k-1}-$ 1 reassignments per job of phase $k$ are necessary under any reassignment strategy. Next we describe the reassignments done by FirstFit. We split the blocks of jobs and the time intervals in half; block $B_{b}$ is thus split in half-blocks $B_{b}^{1}$ and $B_{b}^{2}$ of size $L_{k}=\frac{1}{2}L_{k-1}$, which are assigned to the interval
$I_{b}^{1}\coloneqq\left[L_{k-1}(2^{k-1}-b-1)+1,\ L_{k-1}(2^{k-1}-b-1+\frac{1}{2})\right]$
and
$I_{b}^{2}\coloneqq\left[L_{k}(2^{k}-2b-1)+1,\ L_{k-1}(2^{k-1}-b)\right],$
respectively. Jobs of phase $k$ build a further block $B_{2^{k-1}}=B_{2^{k-1}}$ of size $L_{k}$. Now consider a reassignment sequence which shifts the $i$-th job of each block one block to the right, assigning the job of block $B_{1}$ to the $i$-th empty slot in interval $I_{0}$:
$\varphi\colon j_{(b,i)}\mapsto\alpha_{k-1}(j_{(b-1,i)})\qquad\text{for }i\in\{1,\ldots,L_{k}\}\text{ and }b\in\{1,\ldots,2^{k-1}\}.$
Observe that only the first $L_{k}$ jobs of each block are involved in the reassignments, which corresponds to the half-blocks $B^{1}_{b}$, $b\in\{1,\ldots,2^{k-1}-1\}$. Assignment $\varphi$ is feasible: consider the $i$-th job $j\coloneqq j^{k-1}_{(b,i)}$ of $B_{b}$ for some $i\in\{1,\ldots,L_{k}\}$ and $b\in\{1,\ldots,2^{k-1}\}$. The deadline of job $j=L_{k-1}(b-1)+i$ is
$d_{j}=N+1-\frac{N}{2^{k-1}}(b-1)-i=\frac{N}{2^{k-1}}(2^{k-1}-b)+\left(\frac{N}{2^{k-1}}-i+1\right),$
and the new slot is
$\varphi(j)=\alpha_{k-1}(j_{(b-1,i)})=L_{k-1}(2^{k-1}-(b-1)-1)+i=\frac{N}{2^{k-1}}(2^{k-1}-b)+i\leqslant d_{j},$
since $i\leqslant\frac{N}{2^{k}}=\frac{1}{2}\cdot\frac{N}{2^{k-1}}$. The reassignment path contains $12^{k-1}-$ 1 jobs being reassigned. Moreover, since reassignments are executed one by one for jobs in order of their arrival, the earliest possible slot is selected at each stage in each reassignment sequence. Hence, the described reassignment $\varphi$ is exactly the FirstFit reassignment. The slot assignment $\alpha_{k}$ at the end of phase $k$ has thus the following structure: the set of jobs is split into $12^{k}-$ 1 blocks $\bar{B}_{\ell}$, $\ell\in\{1,\ldots,2^{k}-1\}$, where $\bar{B}_{\ell}=B^{2}_{b}$ if $\ell=2b$ and $\bar{B}_{\ell}=B^{1}_{b}$ if $\ell=2b-1$. Hence we have
$\alpha_{k}\colon\{1,\ldots,N-\frac{N}{2^{k}}\}\to\{1,\ldots,N-\frac{N}{2^{k}}\},\qquad j\mapsto\begin{cases}\varphi(j),&\text{if }j\in B^{1}_{b}\text{ for some }b,\\ \alpha_{k-1}(j),&\text{otherwise}.\end{cases}$
In other words, blocks $\bar{B}_{\ell}$ for even $\ell=2b$ remain assigned to intervals
$\bar{I}_{\ell}=I_{b}^{2}=\left[L_{k}(2^{k}-2b-1)+1,\ L_{k}(2^{k}-2b)\right]=\left[L_{k}(2^{k}-\ell-1)+1,\ L_{k}(2^{k}-\ell)\right].$
For odd values $\ell=2b-1$, $b\leqslant 2^{k-1}$, blocks $\bar{B}_{\ell}=B_{b}^{1}$ are assigned according to assignment $\varphi$ to the intervals
$\bar{I}_{2b-1}=I_{b-1}^{1}$ $=\left[L_{k}(2^{k}-2(b-1)-2)+1,\ L_{k}(2^{k}-2(b-1)-1)\right]$
$=\left[L_{k}(2^{k}-\ell-1)+1,\ L_{k}(2^{k}-\ell)\right].$
Hence, the assignment $\alpha_{k}$ at the end of phase $k$ has the desired structure.
The $12^{k-1}-$ 1 blocks of size $L_{k}$ are reassigned, yielding $\frac{N}{2^{k}}(2^{k-1}-1)$ reassignments in phase $k$. $\triangleleft$
Finally, we compute the total number of reassignments done by FirstFit. As shown in the claim, in phase $k\leqslant n$, $\frac{N}{2^{k}}(2^{k-1}-1)$ reassignments take place. In the last phase $n+1$, the last job with index $N$ and time window $\{1\}$ arrives, and causes $N-1$ reassignments. The total number of reassignments is thus
$\sum_{k=1}^{n}\frac{N}{2^{k}}(2^{k-1}-1)+N-1$ $=\frac{N}{2}\sum_{k=1}^{n}(1-\frac{1}{2^{k-1}})+N-1$
$=\frac{N}{2}(n-1)-\frac{N}{2}\sum_{k=1}^{n}\frac{1}{2^{k-1}}+N-1$
$=\mathcal{O}(N\cdot n)=\mathcal{O}(N\log N).$