### v4

# Interval-Constrained Bipartite Matching over Time

Abstract
Interval-constrained online bipartite matching problem frequently occurs in medical appointment scheduling: unit-time jobs representing patients arrive online and are assigned to a time slot within their given time interval. We consider a variant of this problem where reassignments are allowed and extend it by a notion of current time, which is decoupled from the job arrival events. As jobs appear, the current point in time gradually advances. Jobs that are assigned to the current time unit become processed, which fixes part of the matching and disables these jobs or slots for reassignments in future steps. We refer to these time-dependent restrictions on reassignments as the over-time property.
We show that FirstFit with reassignments according to the shortest augmenting path rule is $\frac{2}{3}$-competitive with respect to the matching cardinality, and that the bound is tight. Interestingly, this bound holds even if the number of reassignments per job is bound by a constant. For the number of reassignments performed by the algorithm, we show that it is in $\Omega(n\log n)$ in the worst case, where $n$ is the number of patients or jobs on the online side. This result is in line with lower bounds for the number of reassignments in online bipartite matching with reassignments [14], and, similarly to this previous work, we also conjecture that this bound should be tight. Known upper bounds like the $\mathcal{O}(n\log^{2}n)$ for online bipartite matching with reassignments by Bernstein, Holm, and Rotenberg [4] do not transfer directly: while our interval constraints simplify the problem, the over-time property restricts the set of possible reassignments. Moreover, we show that the algorithm maintaining the earliest-deadline-first order in the schedule yields maximum matchings, but at a cost of $\Omega(n^{2})$ reassignments in worst case.
Finally, we consider the generalization of the problem in which the set of feasible slots has arbitrary shape. We show that FirstFit retains its competitivity in this case and that no other deterministic algorithm can be more than $\frac{2}{3}$-competitive.
1 Introduction
Consider the process of assigning medical appointments of equal duration to patients. Patients request appointments one by one and expect an immediate answer. The time interval within which patients require an appointment is determined by their medical condition: the urgency of treatment imposes the deadline, and the amount of time needed to prepare the patient for treatment, e.g., additional diagnostics before a surgery, determines the earliest possible appointment day. As further patients arrive, previously scheduled patients with looser deadlines may be postponed in favor of urgent patients. As time passes, a growing part of the schedule is in the past and therefore cannot be changed, whereas appointment slots in the future can still be reassigned to different patients. The goal is to schedule as many patients as possible and minimize changes to the schedule.
In this work, we model appointment scheduling as a variant of online bipartite matching with reassignments. In this problem, vertices of one part of the graph, here called “slots”, are known in advance, and vertices of the other part, here called “jobs”, which represent patients, are revealed one by one together with their incident edges. Upon arrival, each job is either immediately assigned to an adjacent slot or irrevocably rejected. It is possible for the algorithm to reallocate jobs to different slots to make place for a new job. The task is to maximize the number of assigned jobs and minimize the number of reallocations.
We extend this well-studied problem by properties motivated in the practical application. We introduce a notion of time to the online problem. As time passes, more jobs arrive and slots are processed, so that once a slot is in the past, jobs assigned to it cannot be reallocated anymore. Furthermore, our jobs feature interval constraints: a job can be allocated only to a slot between its given earliest and latest slot. We refer to the described problem as Interval-Constrained Bipartite Matching over Time, or ICBMT for short, and formalize the problem in Section 1.3. Then, we analyze two algorithmic approaches with respect to the number of scheduled jobs and the required reallocations.
In the motivating medical application, slots are often assigned in a first-come-first-served manner, which is then revoked in case of emergencies. This corresponds to a FirstFit algorithm that uses shortest augmenting paths (SAP) to augment the schedule when there is no feasible slot available. In addition to practical relevance, this algorithm is a natural choice. Due to the over-time property of the problem, algorithms that leave early slots unoccupied risk missing a possible match or have to augment the matching for the current time slot. We analyze this online algorithm and give a tight bound of $\frac{2}{3}$ on the competitive ratio with respect to the cardinality of the resulting matching. Furthermore, we show that if the algorithm always selects the lexicographically smallest (lex-min) shortest path, then it incurs $\Omega(n\log n)$ reassignments in the worst case. This bound can also be adopted for the lex-max paths in a different worst-case instance.
Additionally, we consider the earliest-deadline-first strategy, which is inspired by an optimal strategy for related machine scheduling problems, and show its optimality with respect to the number of scheduled jobs. We also analyze both algorithms in a more general setting where each job indicates a set of feasible slots that is not necessarily connected, instead of an interval, and show that FirstFit maximizes the number of assigned jobs in this setting.
1.1 Related work
ICBMT is a special case of the online bipartite matching problem and of the online assignment problem. In these problems, the servers (possibly with capacity greater than one) are known in advance, and clients are revealed online, each with a subset of servers the client can be served by. Typical objectives are to maximize the number of assigned clients [18] or to balance the load of the servers [5].
In the classical maximum online bipartite matching, all matching decisions are irrevocable and the aim is to maximize the number of matched clients. No deterministic algorithm can have a competitive ratio greater than $\frac{1}{2}$, which is achieved by a greedy algorithm. Karp, Vazirani, and Vazirani present a randomized algorithm ranking, which is strictly $(1-\frac{1}{e})$-competitive, and show the tightness of the bound [18]. The proof has later been corrected, refined, and generalized [13, 6, 11].
In many applications, it is reasonable to allow reassignments; that is, client rejections are irrevocable, but the assigned server of an already assigned client can be changed. In this problem, it is trivial to achieve a maximum matching at the cost of a large number of reassignments. One widely studied strategy for online bipartite matching with reassignments is Shortest Augmenting Paths (SAP), which is also the basis of an optimal algorithm for offline maximum bipartite matching [17]. Grove et al. show a lower bound of $\Omega(n\log n)$ reassignments for SAP [14]. They also show a matching upper bound of $\mathcal{O}(n\log n)$ reassignments for the case of all clients having a degree of two. This upper bound for SAP is later also proven for arrivals in random order [9] and for trees [8]. For general graphs, Bernstein, Holm, and Rotenberg show an amortized upper bound of $\mathcal{O}(n\log^{2}n)$ reassignments for SAP [4]. Chaudhuri et al. conjecture that the lower bound is tight for general bipartite graphs [9].
Further exact approaches to online bipartite matching with reassignments, which are different from SAP, have also been studied [7, 9] and do not improve the number of reassignments.
Another branch of research concerns approximation algorithms for online bipartite matching. Bosek et al. construct $(1-\epsilon)$-approximate matchings applying $\mathcal{O}((1-\epsilon)n)$ reassignments [7]. Bernstein et al. approximate the problem for every $p$-norm simultaneously. Their algorithm is $8$-competitive with $O(n)$ reassignments [5]. Gupta, Kumar, and Stein consider the assignment version of the problem, where each server can serve multiple clients, and present a $2$-competitive algorithm for maximum load minimization with at most $O(n)$ reassignments [15]. Recently, Shin et al. consider the online bipartite matching problem under a bounded number $k$ of allowed reassignments per arrival [21]. They show that a modified SAP-algorithm is $(1-\frac{2}{k+2})$-competitive for this setting and that this bound is tight already on paths.
The influence of recourse has also been studied for more general matching problems, such as minimum-cost matching [20, 16] and maximum matching on general graphs [1, 3, 19].
Another closely related problem is unit job scheduling with integer release times and deadlines. The interval constraints directly correspond to release times and deadlines, and the maximum cardinality objective corresponds to throughput maximization. Here, jobs can be buffered until they are executed or their deadline arrives. The offline version of this problem is optimally solved by the Earliest Deadline First strategy, which can be implemented in $O(n)$ time [12, 22]. For the online version, Baruah, Haritsa and Sharma show that any algorithm that is never idle is at least 2-competitive [2]. Chrobak et al. give a barely random algorithm, an algorithm that only uses a single random bit, that is $\frac{5}{3}$-competitive [10]. Note that in these scheduling problems, deadlines and release dates are not necessarily integers. Moreover, in contrast to ICBMT, decisions on rejections of jobs do not have to be met immediately upon job’s arrival, as jobs can remain pending for several time steps.
1.2 Our contribution
We consider a FirstFit-type algorithm that uses a shortest augmenting path in case there is no directly available slot. For this algorithm, we give a tight lower and upper bound of $\frac{2}{3}$ on the competitive ratio with respect to the cardinality of the final matching. In contrast to related problems that use recourse, the over-time nature of our problem eliminates augmenting paths that may be needed to recover an offline optimal solution to the problem.
In addition, we give an instance on which the algorithm performs $\Omega(n\log n)$ reassignments. To show this result, we have to specify the algorithm more precisely. For every new assignment, there might be several shortest augmenting paths ending in the same slot, and the choice of one impacts the result. Our algorithm always selects the lexicographically smallest shortest path, where the order is with respect to the sequence of slots in the path. This is a natural extension of the FirstFit strategy, which also favors early slots for the current job. We conjecture that there is no deterministic strategy for selecting the shortest paths that yields strictly better results, but a randomization over different strategies might be promising.
Our results imply that with only a constant number of reassignments for each new assignment, we retain the competitive ratio of $\frac{2}{3}$ for only a linear number of reassignments in total. At first glance, this result sounds amazing, but one should not mistake this for the superior algorithm. For every number $k\in\mathbb{N}$ of reassignments allowed per single new assignment, there exists an instance on which this restricted algorithm only achieves a $\frac{k+1}{k+2}$-fraction of the assignment constructed by the unconstrained FirstFit strategy on the same instance.
All these results are in contrast to the fact that the earliest-deadline-first strategy guarantees a maximum matching; that is, it is $1$-competitive with respect to the cardinality of the final matching. However, the strategy makes $\Omega(n^{2})$ reassignments in the worst case.
Our main results are summarized in Table 1.
FirstFit $k$-FirstFit Online-EDF
assigned jobs $\frac{2}{3}$ $\frac{2}{3}$ $1$
(competitive ratio)
reassignments $\Omega(n\log n)$ $\Theta(n)$ $\Theta(n^{2})$
Table 1: Performance bounds for algorithms presented in this work on instances with $n$ jobs.
Furthermore, we consider a few special cases. If all intervals have the same fixed length, we show that FirstFit maximizes the cardinality of the matching in $\Theta(n)$ reassignments. And if the algorithms are allowed to make decisions in batches for all jobs that arrive in the same time unit, the competitive ratio does not improve for the algorithms that we consider. The number of reassignments in this model is improved at most by a factor of $2$.
Finally, we consider the generalization of our problem setting, in which the sets of feasible slots for each job are not necessarily intervals. For this setting, we show that FirstFit remains $\frac{2}{3}$-competitive, and that it is the best possible competitive ratio any deterministic algorithm can attain.
1.3 Problem definitions
In Interval-Constrained Bipartite Matching over Time, we are given processing capacity of one slot per time unit $t\in\mathbb{N}$; we identify slots with time units they belong to. Jobs $j\in\{1,\ldots,n\}$ are revealed one by one together with their arrival times $a_{j}\in\mathbb{N}_{0}$ and their intervals $[r_{j},d_{j}]$, $r_{j},d_{j}\in\mathbb{N}$, which is the set of their feasible slots. We assume all numbers to be polynomial in the (unknown) number of jobs. The arrival times are non-decreasing in the order of jobs’ arrival, i.e., $a_{j}\leqslant a_{j+1}$ for any $j<n$. Several jobs can have the same arrival time, i.e. arrive withing the same time unit, but not simultaneously; the order in which they arrive is fixed and, upon arrival of the first such job, the remaining jobs with the same arrival time are not yet visible to the algorithm. For all jobs $j\leqslant n$, the preparation time $r_{j}-a_{j}$ is at least $1$. The jobs will be denoted as tuples of their attributes $j=(a_{j},r_{j},d_{j})$, and instances as sequences of jobs.
Upon arrival, job $j$ must be assigned to some slot $t\in[r_{j},d_{j}]$. The already scheduled jobs can be reassigned to different slots to free a slot in the job $j$’s interval, but they cannot be rejected. If no reassignment combination allows to assign job $j$, the job is irrevocably rejected. The assignments in slots $t\leqslant a_{j}$ are fixed and cannot be involved in reassignments for job $j$. This models the fact that the appointment schedule for the days already past, including the current day, cannot be changed. We therefore call the slots up to $a_{j}$ fixed.
The goal is to produce online an assignment $\alpha\colon\{1,\ldots,n\}\to\mathbb{N}\cup\{\bot\}$ of jobs to slots, where $\alpha(j)=\bot$ denotes that the job is rejected. The assignment should optimize the following objectives:
  • maximize the number of assigned jobs,
  • minimize the number of reassignments.
We consider the two objectives separately.
Any instance of ICBMT has an underlying bipartite graph consisting of the job set on one side and all feasible slots on the other, with edges connecting jobs to their feasible slots. An offline augmenting path for a solution $\alpha\colon\mathcal{J}\to\mathbb{N}\cup\{\bot\}$ of ICBMT is an augmenting path with respect to the matching $\alpha$ in the underlying bipartite graph.
Remark 1.
An ICBMT solution has an optimal number of assigned jobs if and only if it admits no offline augmenting paths.
In fact, an optimal offline solution for ICBMT is a maximum matching in the underlying bipartite graph.
The main difference between ICBMT and the classical online bipartite matching with reassignments is the temporal component of ICBMT, which leads to a partially fixed assignment as time progresses. However, if all jobs in an instance arrive at the same time, then ICBMT with the objective of minimizing rejections reduces to the online maximum bipartite matching. We call this special case Static ICBMT problem.
Definition 2.
Static ICBMT is a special case of the ICBMT problem, in which all jobs of an instance have the same arrival time, i.e., for any instance $I=\left((a_{j},r_{j},d_{j})\right)_{j=1}^{n}$ of Static ICBMT holds $a_{j}=a\ \text{for all}\ j\in\{1,\ldots,n\}$.
Any instance of Static ICBMT is also an instance of online bipartite matching with reassignments: Indeed, since the arrival time of jobs is constant, no part of the assignment becomes fixed, and all reassignments are allowed at any time. Hence, a feasible solution for online bipartite matching is also a feasible solution for Static ICBMT on the same set of jobs.
2 Greedy approaches
One natural and widely used appointment assignment strategy is to greedily assign the earliest possible slot to every patient. If necessary, a few other patients are reassigned.
The algorithm we call FirstFit in this work follows this greedy paradigm: Assign jobs to early slots and minimize the number of reassignments for each job arrival. It is a special case of the greedy shortest augmenting path strategy for bipartite matching that additionally considers the special structure of instances of ICBMT, namely the total order on the set of slots. The algorithm is defined as follows.
The input consists of a job $j^{\prime}=(a_{j^{\prime}},r_{j^{\prime}},d_{j^{\prime}})$ and a current slot assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$, where $\mathcal{J}\subseteq\{1,\ldots,n\}$ is the set of previously assigned jobs. FirstFit searches for a shortest sequence of reassignments which frees up one slot in the interval of $j^{\prime}$ and assigns the last job of the sequence to a free slot $t^{*}$. If no such sequence of reassignments exists, job $j^{\prime}$ is rejected.
The sequence of reassignments results from a shortest augmenting path $P$ from job $j^{\prime}$ to a free slot $t^{*}$ in a bipartite job-slot graph $G_{\alpha,j^{\prime}}$. The graph $G_{\alpha,j^{\prime}}=(V^{T}\,\dot{\cup}\,V^{J},E^{\alpha}\,\dot{\cup}\,E^{\overline{\alpha}})$ is a digraph containing a vertex for each non-fixed slot and job: $V^{T}=\left\{t\in\cup_{j\in V^{J}}[r_{j},d_{j}]\mid t>a_{j^{\prime}}\right\}$ and $V^{J}=\left\{j\in\mathcal{J}\mid\alpha(j)>a_{j^{\prime}}\right\}\,\dot{\cup}\,\{j^{\prime}\}$. There are edges $E^{\overline{\alpha}}$ from each job to its feasible slots it is not assigned to, $E^{\overline{\alpha}}=\{(j,t)\mid j\in V^{J},\ t\in V^{T},\ t\in[r_{j},d_{j}],\ \alpha(j)\neq t\}$, and edges $E^{\alpha}$ from each filled slot to the job assigned to it, $E^{\alpha}=\{(t,j)\mid t\in V^{T},\ j\in V^{J},\ \alpha(j)=t\}$. The path $P$ is augmenting with respect to the current assignment $\alpha$. Among multiple possible shortest paths, we choose the one with the earliest target slot, which motivates the name FirstFit. If such a path exists, then the new assignment $\alpha^{\prime}$ is a symmetric difference between the edges of the old assignment and of the path $P$, denoted $\alpha\Delta\{e\mid e\in P\}$.
Input: job $j^{\prime}=(a_{j^{\prime}},r_{j^{\prime}},d_{j^{\prime}})$, current slot assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$
Output: a new assignment $\alpha^{\prime}\colon\mathcal{J}\cup\{j^{\prime}\}\to\mathbb{N}\cup\{\bot\}$
$F\coloneqq\{t\in\mathbb{N}\mid t>a_{j^{\prime}}\,\land\,\alpha^{-1}(t)=\emptyset\}$;
  /* free slots */
$\mathcal{P}\coloneqq\{p\mid p\text{ is a shortest augmenting path from }j^{\prime}\text{ to }F\}$ ;
$P\coloneqq\operatorname*{arg\,min}\{t\in F\mid t\text{ is the endpoint of }p\in\mathcal{P}\}$ ;
  /* earliest target slot */
if $P$ exists then
       $\alpha^{\prime}\coloneqq\alpha\,\Delta\,\left\{(j,t)\mid(j,t)\in P\text{ or }(t,j)\in P\right\}$;
else
       $\alpha^{\prime}\coloneqq\alpha\cup(j^{\prime},\bot)$;
return $\alpha^{\prime}$
Algorithm 1 FirstFit
Note that we do not need to enumerate the set of paths $\mathcal{P}$: the earliest-target shortest augmenting path can be found efficiently via breadth-first search on $G_{\alpha,j^{\prime}}$, if the nodes of one level are processed in the chronological order. Notice that FirstFit can also be used for Bipartite Matching over Time without interval constraints, where sets of feasible slots of jobs are of arbitrary structure (see Section 5). On interval-constrained instances, however, we additionally have the following property.
Remark 3.
When assigning job $j^{\prime}$, FirstFit for ICBMT always chooses the earliest free slot in the assignment in or after the interval of job $j^{\prime}$, i.e., slot $t^{*}\coloneqq\min\{t\geqslant r_{j^{\prime}}\mid\alpha^{-1}(t)=\emptyset\}$, as the target free slot.
If multiple shortest paths to the earliest possible slot exist, one is selected arbitrarily. We will discuss the influence of the choice of the path on the algorithm’s performance at the end of Section 2.
2.1 Matching size under FirstFit
We start our analysis of FirstFit with the primary objective in ICBMT – the number of assigned jobs, i.e., the cardinality of the resulting matching. In general, FirstFit can make unnecessary jobs rejections, as we will see later. On instances of Static ICBMT, however, FirstFit achieves a maximum number of assigned jobs.
Observation 4.
FirstFit yields a maximum matching on instances of Static ICBMT.
In fact, as mentioned above, Static ICBMT is a special case of maximum online bipartite matching. FirstFit, in turn, coincides with the SAP strategy for online bipartite matching complemented with a tie-breaking rule between paths of the same length: The earliest terminal slot is preferred. Hence, the optimality of FirstFit follows from the optimality of SAP [9].
To prove the next optimality result, we introduce the notion of a closed interval.
Definition 5.
A closed interval in an assignment $\alpha\colon\mathcal{J}\to\mathbb{N}\cup\{\bot\}$ for ICBMT is a time interval $I\subseteq\mathbb{N}$ without empty slots such that all jobs $j\in\alpha^{-1}(I)$ have intervals $[r_{j},d_{j}]\subseteq I$.
Observe that any alternating path with respect to the matching $\alpha$ in the underlying bipartite graph that starts in a closed interval can visit only slots of this interval and jobs assigned to them.
Theorem 6.
If all jobs in an instance of ICBMT have intervals of the same length, then FirstFit yields a maximum matching.
Theorem 7.
If jobs have at least two different interval lengths, then FirstFit is at most $\frac{2}{3}$-competitive for ICBMT with respect to the cardinality of the matching.
Proof.
We define two job types: Urgent jobs with preparation time of $1$ and interval length $\delta$, and regular jobs with preparation time $2$ and intervals of length $\Delta\geqslant 3\delta-2$. Consider now the following instance $\mathcal{I}$ of ICBMT (see Figure 1). First, $\ell\coloneqq\delta-1$ regular jobs $j_{i}$ for $i\in\{1,\ldots,\ell\}$ arrive at time $-1$ (we extend the time scale to negative numbers to simplify the notation). Their interval is $[1,\Delta]$, and FirstFit assigns them to slots $\alpha(i)=i$ for $i\in\{1,\ldots,\ell\}$. Next, for each $i\in\{1,\ldots,\delta\}$, an urgent job $u_{i}$ with interval $[i,i+\ell]$ is revealed at time $i-1$. Upon arrival of job $u_{i}$, the first $\ell+i-1$ slots are occupied, so $u_{i}$ is assigned to slot $\ell+i$, the last slot of its interval.
Finally, further $\ell$ urgent jobs $v_{i}$, $i\in\{1,\ldots,\ell\}$ with interval $[\delta,\delta+\ell]$ are revealed also at time $\ell$. All slots in their interval are occupied. The assignment up to the current time point $\ell$ cannot be modified, so the regular jobs cannot be reassigned.
$t$$\ell$$\delta$$\Delta$$\ell+\delta$$\ell=\delta-1$$\delta$$\ell$
Figure 1: Upper bound instance for $\delta=4$, $\ell=3$. The arrival order of jobs is from top to bottom. The arrival times are marked with dots and the FirstFit assignment with yellow circles. The grayed-out part of the solution is fixed upon arrival of the last $\ell$ jobs. The dashed jobs are rejected.
Hence, jobs $v_{i}$, $i\in\{1,\ldots,\ell\}$, are all rejected by FirstFit, so FirstFit achieves the objective value $\text{FF}(\mathcal{I})=2\delta-1$. An optimal assignment $\beta$ assigns all jobs to a feasible slot:
$\begin{array}[]{ll}\beta(j_{i})\,=\delta+\ell+i,&1\leqslant i\leqslant\ell,\\ \beta(u_{i})=i,&1\leqslant i\leqslant\delta,\\ \beta(v_{i})=\delta+i,&1\leqslant i\leqslant\ell;\end{array}$
hence, $\text{OPT}(\mathcal{I})=3\delta-2$. From
$\frac{\text{FF}(\mathcal{I})}{\text{OPT}(\mathcal{I})}=\frac{2\delta-1}{3\delta-2}\underset{\delta\to\infty}{\longrightarrow}\frac{2}{3}$
follows that the competitive ratio of FirstFit is not greater than $\frac{2}{3}$. ∎
We complement this result with a tight lower bound. Even if we drop the interval constraints and allow for arbitrary sets of feasible slots, FirstFit is at least $\frac{2}{3}$-competitive.
Theorem 8.
FirstFit is at least $\frac{2}{3}$-competitive for ICBMT with respect to the number of assigned jobs.
Proof.
Consider an arbitrary instance of ICBMT. Let $\mathcal{J}_{\text{FF}}$ be the set of jobs assigned by FirstFit, and $\mathcal{J}_{\text{OPT}}$ be the set of jobs assigned in an optimal solution; let $f\colon\mathcal{J}_{\text{FF}}\to\mathbb{N}$ and $o\colon\mathcal{J}_{\text{OPT}}\to\mathbb{N}$ be the corresponding assignments.
To compare the cardinalities of the two assignments, we construct alternating job-slot sequences that help to partition the job set. For each job $j_{0}\in\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}$ rejected by FirstFit but assigned by the optimal algorithm, we construct a sequence $Q(j_{0})$ by first adding job $j_{0}$ and its optimal slot $s_{0}=o(j_{0})$ to it. Next, in each iteration $k\in\mathbb{N}$, we consider the job $j_{k}=f^{-1}(s_{k-1})$ that is assigned by FirstFit to the optimal slot $s_{k-1}$ of the previous job in the sequence. Job $j_{k}$ and its optimal slot $s_{k}=o(j_{k})$ are added to the sequence $Q(j_{0})$. We repeat these iterations until we either reach a slot $s_{k}$ that is free in $f$ or a job $j_{k}$ that is not assigned by $o$.
Our construction yields two types of sequences. In each sequence, the first job is in $\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}$, the intermediate jobs are in $\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}$, and the last job is either in $\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}$ as well, if the sequence ends with a slot, or in $\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}$. We call sequences that end with a job in $\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}$ and a slot even, and those that end in $\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}$ odd, according to the length of the sequence.
We show that each even sequence contains at least three jobs. If a sequence $Q(j_{0})$ for some rejected job $j_{0}$ contains only one job, then $Q(j_{0})=(j_{0},o(j_{0}))$, and slot $o(j_{0})$ is free in assignment $f$. So FirstFit would have assigned $j_{0}$ to this slot.
Now consider an even sequence with two jobs $Q(j_{0})=(j_{0},s_{0},j_{1},s_{1})$, where $s_{0}=o(j_{0})$, $s_{1}=o(j_{1})$ and $f(j_{1})=s_{0}$ (see Figure 2).
$j_{0}$$j_{1}$$s_{0}$$s_{1}$
Figure 2: An even alternating sequence with two jobs.
Since slot $s_{1}$ is free in the FirstFit solution and job $j_{1}$ is assigned to slot $s_{0}$, we have $s_{0}<s_{1}$. Furthermore, we argue that $j_{1}$ arrived before $j_{0}$ and that upon arrival of $j_{0}$, job $j_{1}$ was already assigned to $s_{0}$. If this is the case, then the path $(j_{0},s_{0},j_{1},s_{1})$ has been an augmenting path for $j_{0}$, so FirstFit would reassign $j_{1}$ instead of rejecting job $j_{0}$. This augmenting path is feasible because all its edges are contained in the optimal assignment $o$ or in the FirstFit assigment $f$.
So suppose that, upon arrival of job $j_{0}$, slot $s_{0}$ was occupied not by $j_{1}$, but by some further job $j_{2}$. In the final FirstFit assignment we have $f(j_{1})=s_{0}$; hence, at some later time point, job $j_{1}$ is reassigned, or assigned straight upon its arrival, to $s_{0}$. Slot $s_{1}$ is still available at that time; hence, the augmenting path (re-)assigning $j_{1}$ to $s_{0}$ can be shortened by assigning $j_{1}$ to a free slot $s_{1}$. This contradicts slot $s_{1}$ remaining empty in the final FirstFit assignment. So $j_{1}$ was occupying slot $s_{0}$ upon arrival of $j_{0}$, and FirstFit would have assigned $j_{0}$ by reassigning $j_{1}$.
Next, we express the number of jobs assigned or rejected by both algorithms through the number of odd and even sequences $n_{\text{odd}}$ and $n_{\text{even}}$, respectively. For the odd sequences, we use the trivial bound $n_{\text{odd}}\leqslant\left|\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}\right|$. For the even sequences, we have shown that each sequence contains at least three jobs. This yields a bound of $n_{\text{even}}\leqslant\frac{1}{2}\left|\mathcal{J}_{\text{FF}}\cap\mathcal{J}_{\text{OPT}}\right|$. As jobs in $\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}$ only occur at the start of a sequence, we have
$\left|\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}\right|=n_{\text{odd}}+n_{\text{even}}\leqslant\left|\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}\right|+\frac{1}{2}\left|\mathcal{J}_{\text{FF}}\cap\mathcal{J}_{\text{OPT}}\right|\;.$
Finally, we obtain
$\left|\mathcal{J}_{\text{OPT}}\right|$ $=\left|\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}\right|+\left|\mathcal{J}_{\text{OPT}}\setminus\mathcal{J}_{\text{FF}}\right|\leqslant\frac{3}{2}\left|\mathcal{J}_{\text{OPT}}\cap\mathcal{J}_{\text{FF}}\right|+\left|\mathcal{J}_{\text{FF}}\setminus\mathcal{J}_{\text{OPT}}\right|\leqslant\frac{3}{2}\left|\mathcal{J}_{\text{FF}}\right|\;.$
2.2 Reassignments by FirstFit
Another natural performance measure for online algorithms for ICBMT is the number of reassignments made by the algorithm. Before we present our analysis of FirstFit for this measure, note that we did not specify how ties between shortest augmenting paths from the current job to the earliest free slot are resolved. Although each of the shortest paths yields the same number of reassignments in that iteration, we will see that the resulting solutions lead to different numbers of reassignments in later iterations.
All possible shortest augmenting paths can be uniquely represented and compared by the sequence of slots they visit on the way from the new job to the target slot. One possible rule for path selection is to choose the lexicographically smallest sequence of slots. Intuitively, this rule favors early slots for incoming jobs, and the path can be found in polynomial time.
We show in Lemma 9 that with this path choice rule, which we denote lex-min for short, FirstFit makes at least $\Omega(n\log n)$ reassignments in the worst case. Observe that this bound coincides with the lower bound on the worst-case number of reassignments for the SAP strategy for online bipartite matching [14].
Lemma 9.
In the worst case, FirstFit makes $\Omega(n\log n)$ reassignments on instances with $n$ jobs, even on instances of Static ICBMT.
Proof.
Here, we provide only a sketch of the proof; for the full proof, see Appendix A.1.
Consider an upper-triangle instance with $N=2^{n}$ jobs arriving at time 00. Jobs are enumerated in the order of arrival, job $i\in\{1,\ldots,N\}$ has interval $[1,N+1-i]$, see Figure 3.
To describe how FirstFit works on this instance, we divide the set of jobs into $n+1$ subsequent phases: Phase $k\leqslant n$ contains $\frac{N}{2^{k}}$ jobs, and phase $n+1$ contains one job.
In the first phase, the first half of all jobs arrives and is assigned by FirstFit to the first $\frac{N}{2}$ slots. In the second phase, the half of the remaining jobs arrives. As there is no free slot for any of these new jobs, each new job causes a reassignment of one of the jobs of phase one. According to the lex-min rule for choosing augmenting paths, jobs of phase $2$ are assigned to the first $\frac{N}{4}$ slots, and the first $\frac{N}{4}$ jobs of phase $1$ are reassigned to the third block of $\frac{N}{4}$ slots, i.e., to slots $\frac{N}{2}+1,\ldots,\frac{3N}{4}$.
In each subsequent phase $k$, the half of already present jobs gets reassigned. The chain of reassignments caused by job $i$ ends with slot $i$, and each job moves $\frac{N}{2^{k-1}}$ slots to the right, so each chain contains $12^{k-1}-$ 1 reassignments (see Figure 3). Hence, in phase $k$ there are $\frac{N}{2^{k}}(2^{k-1}-1)=\frac{N}{2}-\frac{N}{2^{k}}$ reassignments in total. For detailed calculations, see Appendix A.1.
$t$Phase 1Phase 2Phase 3
Figure 3: Example instance with the reassignments in phase $3$. Under lex-min rule, jobs of the current phase occupy the first slots in their order of arrival, and all other jobs are assigned in parallel diagonal blocks. The reassignment chains propagate this block structure through the instance.
Hence, the total number of reassignments up to phase $\log N$ is
$\sum_{k=2}^{\log N}\left(\frac{N}{2}-\frac{N}{2^{k}}\right)$ $=\frac{N}{2}(\log(N)-1)-N\sum_{k=2}^{\log N}\frac{1}{2^{k}}$
$=\frac{N}{2}(\log(N)-1)-N\left(\frac{1}{2}-\frac{1}{N}\right)\in\Omega(N\log N)\;.$
Observe that on the ICBMT instance in Lemma 9, a different path choice rule, e.g., the lex-max rule that chooses the lexicographically maximal sequence of visited slots, would lead to only $\frac{N}{2}$ reassignments in total.
However, the lex-max rule is also sub-optimal: Consider an instance similar to that in Figure 3, but in which the second half of the jobs have right-aligned intervals $[i-\frac{N}{2},\frac{N}{2}]$ for $i\in\{\frac{N}{2}+1,\ldots,N\}$. By reasoning analogously to the proof of Lemma 9, the lex-max path choice rule yields $\mathcal{O}(n\log n)$ reassignments on this instance, whereas lex-min requires only $\frac{N}{2}$ reassignments.
As for the upper bound on the number of reassignments, there is a known upper bound of $\mathcal{O}(n\log^{2}n)$ reassignments for online bipartite matching under the Shortest Augmenting Path strategy. This bound holds in particular for Static ICBMT as a special case of online bipartite matching, but not necessarily for ICBMT in general. Unfortunately, we were unable to show an analogous upper bound for ICBMT. Our problem differs from online bipartite matching in two aspects: The interval structure of feasible sets and the temporal component that fixes parts of the graph during the online arrival process. While the interval constraints possibly reduce the complexity of the problem by restricting possible adversarial instances, the temporal aspect of ICBMT instances makes any SAP-based algorithm less potent, since there are fewer augmenting paths available, as Figure 4 demonstrates.
$t$$n-4$
Figure 4: An instance with $n=8$ jobs with the FirstFit assignment, identical to an SAP assignment, upon arrival of the last job. The reassignments by FirstFit in ICBMT are shown in red. Online bipartite matching admits a shorter augmenting path of length $2$, shown in blue dashed.
Finally, we again consider a special case of instances with uniform interval length, assuming that these instances allow for better bounds also with respect to the number of reassignments.
Lemma 10.
On instances of ICBMT with fixed uniform interval length, FirstFit makes $\Theta(n)$ reassignments in the worst case.
2.2.1 Limited number of reassignments
Next, we introduce a modification of FirstFit which makes a limited number of reassignments per job arrival, analogously to the modified SAP strategy considered by Shin et al. [21]. FirstFit with path limit $k$, or $k$-FirstFit for short (Alg. 2), is only allowed to use augmenting paths whose length does not exceed $12k+$ 1 for a fixed $k\in\mathbb{N}$, that is, paths which contain at most $k$ already scheduled jobs. Thus, k-FirstFit reassigns at most a constant number of jobs per new assignment.
Input: job $j^{\prime}=(a_{j^{\prime}},r_{j^{\prime}},d_{j^{\prime}})$, current slot assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$
Output: new assignment $\alpha^{\prime}\colon\mathcal{J}\cup\{j^{\prime}\}\to\mathbb{N}\cup\{\bot\}$
$t^{*}\coloneqq\min\{t\geqslant r_{j^{\prime}}\mid\alpha^{-1}(t)=\emptyset\}$;
Find a shortest augmenting $j^{\prime}$-$t^{*}$ path $P$ in $G_{\alpha,j^{\prime}}$;
if $\left|P\right|\leqslant 2k+1$ then
       $\alpha^{\prime}\coloneqq\alpha\,\Delta\,\left\{(j,t)\mid(j,t)\in P\text{ or }(t,j)\in P\right\}$;
else
       $\alpha^{\prime}\coloneqq\alpha\cup(j^{\prime},\bot)$;
return $\alpha^{\prime}$
Algorithm 2 FirstFit with path limit $k$ ($k$-FirstFit)
Clearly, $k$-FirstFit makes at most $kn=\mathcal{O}(n)$ reassignments. Interestingly, both Theorem 8 and the example in Theorem 7 also hold for FirstFit with path limit.
To see why Theorem 8 holds for $k$-FirstFit, recall that the only nontrivial part of the proof is showing that alternating sequences of length $4$ do not exist. The argument is based on two main statements. First, a path encoded by an alternating sequence $(j_{0},s_{0},j_{1},s_{1})$ is an augmenting path for the presumably rejected job $j_{0}$. Since the augmenting path uses only one reassignment, it is feasible for $k$-FirstFit with any $k\in\mathbb{N}$. Second, the non-terminal job $j_{1}$ of the sequence was assigned to slot $s_{0}$ already upon arrival of $j_{0}$, as otherwise the path used to later reassign $j_{1}$ would not be the shortest. This argument clearly holds for $k$-FirstFit as well.
Hence, we obtain an assignment strategy which is bounded in both the number of rejections and reassignments.
Corollary 11.
The $k$-FirstFit assignment strategy for ICBMT is $\frac{2}{3}$-competitive with respect to the number of assigned jobs and makes at most $k\cdot n$ reassignments.
Since the competitive ratios of FirstFit and $k$-FirstFit for any $k\in\mathbb{N}$ are equal, there is no clear trade-off between the number of reassignments and the number of accepted jobs. Nevertheless, with a limit on the number of reassignments per new assignment, we expect to obtain worse solutions on average. For instance, for any $k\in\mathbb{N}$, there are instances on which $(k+1)$-FirstFit yields a maximum matching, while $k$-FirstFit assigns only a $\frac{k+1}{k+2}$-fraction of jobs.
In particular, $k$-FirstFit is not optimal for Static ICBMT. Therefore, although $1$-FirstFit has the best worst-case guarantee on the number of reassignments, choosing a bigger $k$ or using unlimited FirstFit might still be preferable.
3 Proactive reassignments
So far, we have assumed that reassignments are made only if it is otherwise impossible to assign a newly arriving job. Now we investigate an assignment algorithm that strategically reassigns jobs upon arrival of new requests.
The Online-EDF algorithm (Alg. 3) mimics the optimal offline assignment strategy Earliest Deadline First (EDF). It maintains a maximal matching in the future part of the schedule and rejects jobs that cannot be matched in addition to the already scheduled ones. Ties in the ordering of jobs in the schedule are broken according to non-decreasing deadlines.
Input: job $j^{\prime}=(a_{j^{\prime}},r_{j^{\prime}},d_{j^{\prime}})$, current slot assignment $\alpha\colon\mathcal{J}\to\mathbb{N}$
Output: new assignment $\alpha^{\prime}\colon\mathcal{J}\cup\{j^{\prime}\}\to\mathbb{N}\cup\{\bot\}$
$S\coloneqq\{j^{\prime}\}\cup\{j\in\mathcal{J}\mid\alpha(j)>a_{j^{\prime}}\text{ and }d_{j}>d_{j^{\prime}}\}$;
sort $S$ by deadlines;
initialize $\alpha^{\prime}\colon j\mapsto\begin{cases}\bot,&j\in S,\\ \alpha(j),&j\in\mathcal{J}\setminus S;\end{cases}$
foreach $j\in S$ do
       if no free slot in $[r_{j},d_{j}]$ then
             $\alpha^{\prime}\coloneqq\alpha\cup(j^{\prime},\bot)$;
              /* reject $j^{\prime}$, keep old assignment */
             return $\alpha^{\prime}$;
      else
             $\alpha^{\prime}(j)\coloneqq\min\{t\in[r_{j},d_{j}]\mid(\alpha^{\prime})^{-1}(t)=\emptyset\}$;
              /* choose earliest free slot */
return $\alpha^{\prime}$
Algorithm 3 Online-EDF
Observe that, in contrast to scheduling problems, where jobs with earlier deadlines are prioritized and may push out other jobs from the schedule, Online-EDF for ICBMT respects the arrival order of jobs and rejects later jobs with an earlier deadline, if they cannot be assigned without discarding some already accepted jobs.
Theorem 12.
Online-EDF maximizes the number of assigned jobs.
Clearly, Online-EDF is inefficient in terms of the number of reassignments. We show that it achieves the trivial upper bound of $\Omega(n^{2})$ reassignments per $n$ jobs.
Lemma 13.
In the worst case, Online-EDF makes $\Omega(n^{2})$ reassignments, even on instances of Static ICBMT.
Proof.
Consider again an instance consisting of $n$ jobs with monotonously decreasing deadlines, as in the proof of Lemma 8 (see Figure 3). Upon each job arrival, all already present jobs will be reassigned one slot further, in order to assign the new job to the first slot. Hence, the algorithm makes $\sum_{i=1}^{n}(i-1)=\frac{n(n-1)}{2}$ reassignments in total. ∎
To complete the comparison of FirstFit and Online-EDF, we consider the number of reassignments of Online-EDF on instances with uniform interval lengths.
Lemma 14.
If all $n$ jobs have the same interval length $\delta$, then Online-EDF makes $\Theta(\delta n)$ reassignments.
Finally, we consider a version of Online-EDF with a limited number of allowed reassignments per iteration, analogously to $k$-FirstFit: $k$-Online-EDF works identically to Online-EDF if an arriving job causes at most $k$ reassignments, and rejects the job otherwise. In contrast to FirstFit, limiting the number of reassignments of Online-EDF reduces its competitivity. Consider instances as in Figure 3, where the $i$-th job in the order of arrival has interval $[1,\ n-i+1]$ for $i\in\{1,\ldots,n\}$ and causes reassignments of all previously accepted jobs. Therefore, $k$-Online-EDF accepts only $k+1$ first jobs and rejects the $n-k-1$ remaining ones. Hence, for a fixed $k\in\mathbb{N}$, $k$-Online-EDF has no constant competitivity ratio.
4 Batching per time unit
In the problem definition in the introduction, we assume that jobs are processed strictly in the order of their arrival. Now we define an extended version of ICBMT, in which we are allowed to buffer the jobs arriving during one time unit and to assign them all at once, thus obtaining potentially better solutions.
Definition 15 (ICBMT with Batching, ICBMT-B).
In ICBMT with Batching, the assignment is recomputed once per time unit. That is, all jobs with the same arrival time are buffered and assigned simultaneously.
We call algorithms for ICBMT-B batching algorithms. They additionally feature a deterministic behavior on input consisting of batches, i.e., sets of jobs. We say that a non-batching algorithm $\mathcal{A}$ and a batching algorithm $\mathcal{A}^{\prime}$ correspond to each other, if their behavior on single jobs is identical.
Lemma 16.
Let $\mathcal{A}$ be an algorithm for ICBMT with a competitive ratio smaller than $c$ for $c\in\mathbb{Q}$. Let $\mathcal{A}^{\prime}$ be the corresponding batching algorithm. Then $\mathcal{A}^{\prime}$ is at most $\frac{c+1}{2}$-competitive.
Proof.
Consider a worst-case instance $\mathcal{I}$ with $n$ jobs on which the algorithm $\mathcal{A}$ yields a matching of cardinality $c\cdot n$. We show that there exists an instance $\mathcal{I}^{\prime}$ with less than $2n$ jobs such that algorithm $\mathcal{A}^{\prime}$ makes the same number of rejections and reassignments on instance $\mathcal{I}$ as algorithm $\mathcal{A}$ on instance $\mathcal{I}$. To this end, we transform the instance $\mathcal{I}$ into $\mathcal{I}^{\prime}$ as follows. For each time unit $t$ with $k>1$ arrivals, we split slot $t$ to $k$ consequent slots $t^{(i)}$, $i\in\{1,\ldots,k\}$. The jobs arriving at time $t$ now receive pairwise different arrival times $t^{(i)}$ in the order of their arrival. Next, we add $k-1$ dummy jobs $\left(-1,t^{(2)},t^{(k)}\right)$ to the instance. The total number of dummy jobs is less than $n$. Due to their intervals, they block all the additional slots created in the instance transformation, so that starting from time 00, only original slots $t^{(1)}$ for $t\in\mathbb{N}$ and original jobs remain in the instance. As the original jobs have pairwise different arrival times, algorithm $\mathcal{A}^{\prime}$ processes them identically to the corresponding non-batching algorithm $\mathcal{A}$. ∎
Hence, batching algorithms make amortized at least the half of the rejections and reassignments per job arrival compared to their counterparts without batching. For the algorithms considered in this work, however, we show a stronger result: Batching does not improve their competitive ratio.
Lemma 17.
The competitive ratios of Online-EDF and FirstFit with respect to the matching size cannot be improved by batching.
Proof.
Online-EDF is already optimal. As for FirstFit, we slightly modify the worst-case instance family described in Theorem 7 (see Figure 1): We make the arrival times of long jobs pairwise different, and force thereby FirstFit with batching to behave identically to FirstFit without batching. Formally, for some $\delta\in\mathbb{N}$ the instance contains the following jobs, indicated as tuples $(a_{j},r_{j},d_{j})$:
[…]
On this instance, FirstFit with batching behaves as FirstFit and rejects $\delta-1$ of $23\delta-$ 2 jobs. ∎
Moreover, Lemma 16 shows that batching has no effect on our asymptotic results on numbers of reassignments.
Corollary 18.
Batching does not improve the worst-case asymptotic number of reassignments of any algorithm for ICBMT.
5 Relaxing the interval constraint
A practical application – scheduling of medical treatments – motivated the interval constraints in ICBMT. Now we consider a generalization of ICBMT in which we relax the interval constraints. In Bipartite Matching over Time (BMT), each job $j$ has an arrival time $a_{j}$ and a finite set $S_{j}\subset\mathbb{N}_{>a_{j}}$ of feasible slots lying after the arrival time of a job. Note that we still allow for reassignments, so our results are independent of the ranking-algorithm due to Karp, Vazirani and Vazirani [18]. At the same time, our reassignments are restricted by the over-time property, in contrast to unconstrained reassignments considered by Bernstein, Holm and Rotenberg [4].
Both FirstFit and Online-EDF are well-defined in the generalized BMT setting as well. For Online-EDF, the latest slot of the set $S_{j}$ is defined to be the deadline $d_{j}$ of a job $j$. In the following, we discuss the competitivity of both algorithms for BMT.
Let us start with Online-EDF, which is optimal for the interval-constrained version of the problem. Without interval constraints, Online-EDF in our formulation given in Algorithm 3 is not better than a greedy algorithm for online bipartite matching.
Lemma 19.
Online-EDF for BMT is at most $\frac{1}{2}$-competitive with respect to the matching cardinality.
Proof.
Consider an arbitrary instance of online bipartite matching consisting of $n$ jobs arriving online and of $n$ slots; enumerate the slots from $1$ to $n$. We transform this instance into an instance of BMT: we add a slot $n+1$ and add it to each job’s set of slots. Finally, we add an additional job $j_{n+1}$ with $S_{j_{n+1}}=\{n+1\}$, which arrives before all other jobs. As all jobs now have the same deadline, Online-EDF sorts the jobs only in the order of arrival. That is, the algorithms makes no reassignments on this instance, since, upon each new job arrival, all jobs that are already present is the schedule have the same deadline and are not considered for the set $S$ of potential reassignments. Hence, Online-EDF is equivalent to a greedy algorithm without reassignments and is at most $\frac{1}{2}$-competitive on this class of instances. ∎
Next, we consider the FirstFit algorithm. The proof of $\frac{2}{3}$-competitivity in Theorem 8 holds for BMT as well, since the proof does not rely on the interval constraints.
Corollary 20.
FirstFit is $\frac{2}{3}$-competitive for BMT with respect to the matching cardinality.
To show that this competitive ratio is tight, we prove an even stronger result.
Theorem 21.
No deterministic algorithm for BMT has a competitive ratio greater than $\frac{2}{3}$.
Proof.
Consider a deterministic algorithm on a following instance of BMT with three jobs in total. First, jobs $1$ and $2$ with slot sets $S_{1}=\{1,3\}$ and $S_{2}=\{1,2\}$ arrive at time $a_{1}=a_{2}=0$ in that order; see Figure 5.
$t$$t$$t$
Figure 5: Worst-case instance of BMT: the figure shows all three possible assignments of the first two jobs and the corresponding choice of the third, dashed job. The third job will be rejected.
Suppose that both jobs are assigned by the algorithm; if not, a competitive ratio greater than $\frac{2}{3}$ cannot be obtained. The third job arrives at time $a_{3}=1$. With its arrival, the assignment of the first two jobs becomes fixed. Job $3$ has a singe feasible slot that is chosen by an adversary depending on the assignment of the first two jobs: if slot $2$ is occupied, then $S_{3}=\{2\}$; otherwise, slot $3$ must be assigned to job $1$, so we set $S_{3}=\{3\}$. In both cases, the third job will be rejected.
By repeating the described block of three jobs multiple times one obtains instances of any desired size for which one third of the jobs will be rejected. ∎
Hence, FirstFit has the best possible competitive ratio for BMT. In Theorem 6 we saw that FirstFit is optimal for ICBMT if the intervals of jobs are uniform. This is not true any more for BMT.
Remark 22.
FirstFit is at most $\frac{2}{3}$-competitive on instances of BMT if all jobs have the same number of feasible slots.
Proof.
We construct a family of instances by modifying the example in Theorem 7 as shown in Figure 6. For an integer $\delta\geqslant 2$, the instance contains $23\delta-$ 2 jobs. First, $\delta-1$ jobs $j_{i}$ for $i\in\{1,\ldots,\delta-1\}$ arrive. Job $j_{i}$ has arrival time $a_{j_{i}}=-1$ and feasible slots $S_{j_{i}}\coloneqq\{i\}\cup[2\delta,3\delta-2]$. The remaining jobs are as in Theorem 7: the are jobs $u_{i}$ with arrival time $i-1$ and feasible slots $S_{u_{i}}\coloneqq[i,i+\delta-1]$ for each $i\in\{1,\ldots,\delta\}$, and $\delta-1$ identical jobs $v_{i}$ with $a_{v_{i}}=\delta-1$ and $S_{v_{i}}\coloneqq[\delta,2\delta-1]$. Each job $j$ has $\left|S_{j}\right|=\delta$ feasible slots.
Analogously the the proof of Theorem 7, the last $\delta-1$ jobs of this instance are rejected by FirstFit, which yields the upper bound of $\frac{2}{3}$ on the competitive ratio.
[…]
Figure 6: An instance of BMT with each job having $\delta=4$ feasible slots. Each line represents one job; the arrival order is from top to bottom. The arrival times are marked with dots and the FirstFit assignment with yellow circles. The grayed-out part of the solution is fixed upon arrival of the last $\delta$ jobs. The dashed jobs are rejected.
Finally, let us discuss the number of reassignments. The lower bounds on the worst-case number of reassignments for ICBMT clearly transfer to the more general BMT. Hence, also on instances of BMT FirstFit re-assigns $\Omega(n\log n)$ and Online-EDF $\Omega(n^{2})$ jobs. The upper bound of $\mathcal{O}(n\log^{2}n)$ reassignments for the SAP strategy due to Bernstein, Holm and Rotenberg [4] might, as mentioned in Section 2, also apply to our over-time setting, i.e., to FirstFit for BMT. For $k$-FirstFit, the linear bound in the number of reassignments obviously holds in the non-interval setting. Just as for ICBMT, this restricted version still remains $\frac{2}{3}$-competitive, due to the arguments presented in Section 2.2.1. Hence, $k$-FirstFit is an optimal algorithm for BMT that makes at most a linear number of reassignments.
Appendix A Omitted proofs
A.1 Proof of Lemma 9
Lemma 9.
In the worst case, FirstFit makes $\Omega(n\log n)$ reassignments on instances with $n$ jobs, even on instances of Static ICBMT.
Proof.
Consider an upper-triangle instance with $N=2^{n}$ jobs arriving at time 00. Jobs are enumerated in the order of arrival, job $i\in\{1,\ldots,N\}$ has time window $[1,N+1-i]$, see Fig. 3.
We consider only the time interval $[1,N]$, since all job time windows are included in this interval. We partition the set of jobs into $n+1$ phases: phase $k\leqslant n$ contains $\frac{N}{2^{k}}=2^{n-k}$ jobs with indices $N-\frac{N}{2^{k-1}}+1,\ \ldots,\ N-\frac{N}{2^{k}}$, and phase $n+1$ contains one last job.
Claim: When processing jobs of phase $k$ for $k\leqslant n$, FirstFit makes $\frac{N}{2^{k}}(2^{k-1}-1)$ reassignments, and at the end of phase $k$, the assignment $\alpha_{k}$ has the following structure. The jobs that have arrived so far are divided into $12^{k}-$ 1 blocks of $\frac{N}{2^{k}}$ subsequent jobs each; block $b\in\mathbb{N}$ contains jobs $\frac{N}{2^{k}}(b-1)+1,\ \ldots,\ \frac{N}{2^{k}}b$. The time horizon $[1,N]$ of the instance is also divided into $2^{k}$ intervals of length $\frac{N}{2^{k}}$. Let $L_{k}\coloneqq\frac{N}{2^{k}}$ denote the block size at the end of phase $k$. The jobs of block $b\leqslant 2^{k}-1$ are assigned to the $(b+1)$-th last interval in the order of arrival, i.e.,
$\alpha_{k}\colon j_{(b,i)}^{k}\mapsto L_{k}(2^{k}-b-1)+i,$
where $j_{(b,i)}^{k}=L_{k}(b-1)+i$ is the $i$-th job of the $b$-th block in phase $k$, $11\leqslant j\leqslant 2^{k}-$ 1, $1\leqslant i\leqslant L_{k}$.
Proof: Proof by induction.
$k=1$: Clearly, the first $\frac{N}{2}$ jobs are assigned by FirstFit to the slots $[1,\frac{N}{2}]$ in the order of their arrival, i.e.,
$\alpha_{1}\colon[1,\frac{N}{2}]\to[1,\frac{N}{2}],\quad i\mapsto i.$
This corresponds to one block of jobs, assigned to the second last (that is, first) interval of length $\frac{N}{2}$. No reassignments are necessary.
$k-1\to k$: Assume that upon arrival of jobs of phase $k$, the slot assignment $\alpha_{k-1}$ has the structure described above; in particular, jobs of block
$B_{b}\coloneqq\left\{L_{k-1}(b-1)+1,\ldots,L_{k-1}(b-1)\right\}$
are assigned to the interval
$I_{b}\coloneqq\left[L_{k-1}(2^{k-1}-b-1)+1,\ L_{k-1}(2^{k-1}-b)\right],$
where $b\in\{1,\ldots,2^{k-1}-1\}$. We denote by $I_{0}\coloneqq[L_{k-1}(2^{k-1}-1),N]$ the rightmost interval, which contains free slots. Observe that the intervals are enumerated from right to left.
First, we show that for any job of phase $k$, at least $12^{k-1}-$ 1 reassignments are necessary. Jobs of phase $k$ have deadlines $\frac{N}{2^{k-1}}$ or less. Hence, they require slots within the leftmost interval $I_{2^{k-1}-1}$, which is completely occupied by jobs of block $B_{2^{k-1}-1}$.
For any $b\in\{1,\ldots,2^{k-1}-1\}$, jobs of block $B_{b}$ have a deadline not greater than the deadline of job $j_{(b,1)}$, which is
$N+1-L_{k-1}(b-1)-1=L_{k-1}(2^{k-1}-b+1),$
which is the right end point of the interval $I_{b-1}$. Hence, for any $b\in\{1,\ldots,2^{k-1}-1\}$, jobs of block $B_{b}$ can be reassigned to interval $I_{b-1}$ the latest, replacing jobs of block $B_{b-1}$. Consequently, any reassignment path from the interval $I_{2^{k-1}-1}$, where the time windows of new jobs are located, to the empty slots in the interval $I_{0}$ will pass through all intermediate blocks and intervals. That is, at least $12^{k-1}-$ 1 reassignments per job of phase $k$ are necessary under any reassignment strategy.
Next we describe the reassignments done by FirstFit. We split the blocks of jobs and the time intervals in half; block $B_{b}$ is thus split in half-blocks $B_{b}^{1}$ and $B_{b}^{2}$ of size $L_{k}=\frac{1}{2}L_{k-1}$, which are assigned to the interval
$I_{b}^{1}\coloneqq\left[L_{k-1}(2^{k-1}-b-1)+1,\ L_{k-1}(2^{k-1}-b-1+\frac{1}{2})\right]$
and
$I_{b}^{2}\coloneqq\left[L_{k}(2^{k}-2b-1)+1,\ L_{k-1}(2^{k-1}-b)\right],$
respectively. The jobs in phase $k$ build a block $B_{2^{k-1}}=B_{2^{k-1}}$ of size $L_{k}$.
Now consider a reassignment sequence which shifts the $i$-th job of each block one block to the right, assigning the job of block $B_{1}$ to the $i$-th empty slot in interval $I_{0}$:
$\varphi\colon j_{(b,i)}\mapsto\alpha_{k-1}(j_{(b-1,i)})\qquad\text{for }i\in\{1,\ldots,L_{k}\}\text{ and }b\in\{1,\ldots,2^{k-1}\}.$
Observe that only the first $L_{k}$ jobs of each block are involved in the reassignments, which corresponds to the half-blocks $B^{1}_{b}$, $b\in\{1,\ldots,2^{k-1}-1\}$.
Assignment $\varphi$ is feasible: consider the $i$-th job $j\coloneqq j^{k-1}_{(b,i)}$ of $B_{b}$ for some $i\in\{1,\ldots,L_{k}\}$ and $b\in\{1,\ldots,2^{k-1}\}$. The deadline of job $j=L_{k-1}(b-1)+i$ is
$d_{j}=N+1-\frac{N}{2^{k-1}}(b-1)-i=\frac{N}{2^{k-1}}(2^{k-1}-b)+\left(\frac{N}{2^{k-1}}-i+1\right),$
and the new slot is
$\varphi(j)=\alpha_{k-1}(j_{(b-1,i)})=L_{k-1}(2^{k-1}-(b-1)-1)+i=\frac{N}{2^{k-1}}(2^{k-1}-b)+i\leqslant d_{j},$
since $i\leqslant\frac{N}{2^{k}}=\frac{1}{2}\cdot\frac{N}{2^{k-1}}$.
The reassignment path contains $12^{k-1}-$ 1 jobs being reassigned. Moreover, since reassignments are executed one by one for jobs in order of their arrival, the earliest possible slot is selected at each stage in each reassignment sequence. Hence, the described reassignment $\varphi$ is exactly the FirstFit reassignment.
The slot assignment $\alpha_{k}$ at the end of phase $k$ thus has the following structure: the set of job s is divided into $12^{k}-$ 1 blocks $\bar{B}_{\ell}$, $\ell\in\{1,\ldots,2^{k}-1\}$, where $\bar{B}_{\ell}=B^{2}_{b}$ if $\ell=2b$ and $\bar{B}_{\ell}=B^{1}_{b}$ if $\ell=2b-1$. Hence we have
$\alpha_{k}\colon\{1,\ldots,N-\frac{N}{2^{k}}\}\to\{1,\ldots,N-\frac{N}{2^{k}}\},\qquad j\mapsto\begin{cases}\varphi(j),&\text{if }j\in B^{1}_{b}\text{ for some }b,\\ \alpha_{k-1}(j),&\text{otherwise}.\end{cases}$
In other words, blocks $\bar{B}_{\ell}$ for even $\ell=2b$ remain assigned to intervals
$\bar{I}_{\ell}=I_{b}^{2}=\left[L_{k}(2^{k}-2b-1)+1,\ L_{k}(2^{k}-2b)\right]=\left[L_{k}(2^{k}-\ell-1)+1,\ L_{k}(2^{k}-\ell)\right].$
For odd values $\ell=2b-1$, $b\leqslant 2^{k-1}$, blocks $\bar{B}_{\ell}=B_{b}^{1}$ are assigned according to assignment $\varphi$ to the intervals
$\bar{I}_{2b-1}=I_{b-1}^{1}$ $=\left[L_{k}(2^{k}-2(b-1)-2)+1,\ L_{k}(2^{k}-2(b-1)-1)\right]$
$=\left[L_{k}(2^{k}-\ell-1)+1,\ L_{k}(2^{k}-\ell)\right].$
Hence, the assignment $\alpha_{k}$ at the end of phase $k$ has the desired structure.
[…]
$\sum_{k=1}^{n}\frac{N}{2^{k}}(2^{k-1}-1)+N-1$ $=\frac{N}{2}\sum_{k=1}^{n}(1-\frac{1}{2^{k-1}})+N-1$
$=\frac{N}{2}(n-1)-\frac{N}{2}\sum_{k=1}^{n}\frac{1}{2^{k-1}}+N-1$
$=\mathcal{O}(N\cdot n)=\mathcal{O}(N\log N).$